### v1

# Coupled Laplacian Eigenmaps for Locally-Aware 3D Rigid Point Cloud Matching

Abstract
Point cloud matching, a crucial technique in computer vision, medical and robotics fields, is primarily concerned with finding correspondences between pairs of point clouds or voxels. In some practical scenarios, emphasizing local differences is crucial for accurately identifying a correct match, thereby enhancing the overall robustness and reliability of the matching process. Commonly used shape descriptors have several limitations and often fail to provide meaningful local insights on the paired geometries. In this work, we propose a new technique, based on graph Laplacian eigenmaps, to match point clouds by taking into account fine local structures. To deal with the order and sign ambiguity of Laplacian eigenmaps, we introduce a new operator, called Coupled Laplacian, that allows to easily generate aligned eigenspaces for multiple rigidly-registered geometries. We show that the similarity between those aligned high-dimensional spaces provides a locally meaningful score to match shapes. We initially evaluate the performance of the proposed technique in a point-wise manner, specifically focusing on the task of object anomaly localization using the MVTec 3D-AD dataset. Additionally, we define a new medical task, called automatic Bone Side Estimation (BSE), which we address through a global similarity score derived from coupled eigenspaces. In order to test it, we propose a benchmark collecting bone surface structures from various public datasets. Our matching technique, based on Coupled Laplacian, outperforms other methods by reaching an impressive accuracy on both tasks. The code to reproduce our experiments is publicly available at https://github.com/matteo-bastico/CoupledLaplacian and in the Supplementary Code.
1 Introduction
Point cloud matching, or more generally 3D shape matching, is a fundamental task in computer vision. It involves finding the closest matching geometry to a target shape within a set of reference shapes [65]. In addition, if the task involves finding rigid transformations that best align the target shape with the reference, it is often part of a registration process. In particular, point-set rigid registration determines the relative transformation needed to align two point clouds without altering their internal structures [41]. This problem is essential for many practical computer vision tasks, such as medical image analysis [3, 61, 33, 48], intelligent vehicles [22, 35], human pose estimation [23] and objects retrieval and tracking [64, 46]. Traditional [8, 20, 57] and probabilistic registration and matching methods [16, 21, 31, 45], while robust, often struggle to optimally align complex geometries, especially in cases with intricate local structures or slight deformations.
Over the years, several methods have been proposed to tackle the challenge of accurate and efficient 3D shape matching and retrieval [9, 5, 11, 53, 51, 65, 73]. Data-driven 3D shape descriptors [54], capturing underlying properties of the shapes under study, are the common denominator of early shape matching techniques. Global descriptors, such as volume and areas descriptors [78], describe the entirety of the shape in one compact representation, often failing to capture local fine details of complex geometries. On the other hand, local descriptors [39, 57] aim to tackle this issue but they generally are sensitive to noise, based on landmarks and they might not capture semantic information [63]. More recently, deep-learned shape descriptors [72, 6] and deep neural networks for shape matching, based on auto-encoders [73] or transformers [59, 68], have been proposed. Despite their good performances, these methods require a huge amount of annotated data for training, which are hard to collect in fields such as medical imaging [38]. Furthermore, non-rigid point cloud matching and retrieval methods [36, 37, 71] are designed to handle shape deformations and, therefore, they might be excessively flexible ignoring fine local details that are not due to deformations, such as anomalies.
In this study, we introduce a novel method for rigid 3D point cloud matching, based on spectral Laplacian eigenmaps [4, 24], which focus on local details. This technique overcomes the previously mentioned limitations and constraints associated to global and local descriptors and, although it can be seen as a linear graph neural network, it does not require training. One of the main reasons that combinatorial and geometric Laplacians are often considered for spectral graph processing is that their eigenvectors possess properties like the classical Fourier basis functions [79]. We exploit these properties to perform locally-aware shape comparison of point clouds, equipped with k-nearest neighbor graph, without relying on additional descriptors. Nevertheless, eigenspaces alignment, including both eigenvalue ordering and sign disambiguity of the eigendecomposition, is required to correctly match shapes in the spectral space. Current State-of-The-Art methods for such alignment are based on Laplacian eigenfunctions matching [42, 60], i.e. matching the histograms of the eigenvectors. The latter is not robust and frequently fails for highly symmetric geometries, such as bones, with small eigenvalue separation. In this work, we introduce a novel operator, called Coupled Laplacian, designed to simultaneously produce aligned eigenspaces for multiple registered geometries. Namely, we show that when two or more shapes, each with their corresponding graphs, are merged into a single graph using artificial cross-edges, the eigendecompositon of the Laplacian derived from such a combined graph yields aligned spectral spaces for each individual component. Furthermore, this method is order-invariant, count-agnostic and landmarks-free, enabling it to handle point clouds without being influenced by their specific arrangement or the number of points contained. Finally, we utilize the distance between these aligned higher-dimensional spectral spaces, which accounts for intricate local structures, as a global or point-wise score for different applications of 3D shape matching, as shown in Fig. 1.
Figure 1: Overview of the proposed technique applied to different tasks. The global similarity between aligned eigenmaps of multiple geometries, generated from the Coupled Laplacian, is used to match bones and predict their body side. While local similarity is exploited for accurate 3D anomaly detection.
Hence, the identification of local differences between two shapes, using the proposed matching technique, allows us to perform 3D object anomaly detection in the spectral domain with high accuracy. We tested the performances of our method on the MVTec 3D-AD dataset [7], recently proposed for unsupervised point cloud industrial anomaly detection and localization. Furthermore, in the context of medical imaging, correctly identifying the side of a bone (left or right side) is crucial for diagnosis, treatment planning, or skeletal analysis [12]. We refer to the task of determining the side of a Bone or anatomical structure as laterality or Side Estimation (BSE). Recent studies on bone landmark detection [19] and statistical shape modeling [32] have highlighted the limitations of manual side identification for ensuring proper functioning. To the best of our knowledge, manual bone markings [2] are currently the only technique used for this purpose. Consequently, automatic laterality determination arises as an interesting, yet challenging, task to assist the development of fully automated pipelines for medical image analysis, such as patient-specific preoperative planning for Total Knee Arthroplasty (TKA) [34] or Anterior Cruciate Ligament Reconstruction (ACLR) [18, 44]. Thanks to the bilateral symmetry of the animal and human body [28, 67], i.e. right and left sides are mirror shapes of one another, the BSE can be defined as a non-trivial shape matching problem. Given a reference, or source, of known side, to assess the side of a target bone, it is enough to find the best match between the source-target pair and its mirrored counterpart, or vice versa. The complexity of this task is a consequence of subtle local differences present in mirrored bones. This makes it an ideal application of the proposed shape matching technique that utilizes Coupled Laplacian eigenmaps as local geometry descriptors. We generate a benchmark for human BSE by extracting surface point clouds of different bones, e.g. femur, hip and tibia, from public datasets [19, 32, 47]. We also discuss, in the Supplementary Material, the cross-species generalization of the proposed matching algorithm from human to animal, and vice-versa, on an internal dataset [40].
Our contributions are summarized as follows:
  • We propose a new method to preform locally-aware 3D rigid point cloud matching, considering fine local structures, based on graph Laplacian eigenmaps.
  • We define the Coupled Laplacian operator for aligned graphs to tackle the order and sign ambiguity issue of the eigendecomposition.
  • Based on the proposed technique, we introduce a new method for automatic BSE to assist fully automated medical pipelines and we propose a benchmark to test it.
  • We extensively evaluate our method on two tasks: anomaly detection and BSE, outperforming previously proposed techniques.
2 Related Works
Early efforts in developing hand-crafted 3D local features for point cloud matching, registration, recognition and retrieval have typically drawn inspiration from 2D descriptors. Several methods, such as Signature of Histograms of OrienTations (SHOT) [58, 66], Rotational Projection Statistics (RoSP) [26] and Unique Shape Context (USC) [66] rely on the estimation of an unique Local Reference Frame (LRF). The latter is usually based on the eigendecomposition of the covariance matrix of the neighbours of a point of interest, which are then projected into the LRF to analyze their geometric properties. For instance, SHOT [58] captures local shape and orientation information by computing histograms of surface normals and point distribution around keypoints. In contrast, LRF-free approaches [10, 56, 57] try to rely just on features that are intrinsically invariant. The most common LRF-free features are Fast Point Feature Hisograms (FPFH) [57] which, similarly to SHOT, calculates histograms of surface normals considering their relationships in a local region around a keypoint. Despite the progress made with hand-crafted 3D local features, they encounter difficulties when it comes to deal with issues like point cloud resolutions, noisy data, and occlusions [27].
On the contrary, spectral-based point cloud descriptors [1, 29, 52, 70, 62, 55, 69] are a category of feature extraction methods that leverage spectral analysis techniques from graph theory and linear algebra to capture the underlying structure and intrinsic geometric properties of point clouds. Among them, shape-DNA [52] is a surface and solid descriptor based on eigenvalue analysis of the Laplace-Beltrami operator. They propose to use the sequence of eigenvalues (spectrum) of the Laplace operator as a fingerprint characterizing the intrinsic geometry of 3D shapes represented as point clouds. This method has been used for shape retrieval, classification, and correspondence. Weinmann et al. [70] also proposed to extract a features set consisting of $8$ eigenvalues-based indices for each 3D point of a cloud. Furthermore, Heat Kernel Signature (HKS) [62] and Wave Kernel Signatures (WKS) [1] are spectral-based descriptors measuring how heat and wave propagate across a shape, having the eigendecomposition as leading element of the computation. Scaled eigenvectors evaluated at each point are instead directly exploited by the Global Point Signature (GPS) [55] to represent a point cloud as a set of infinite-dimensional vectors, characterizing each point within the global context of the surface it belongs. Nevertheless, the vast majority of these works assumes that the eigenvalues of a shape are distinct and, therefore, can be ordered. Indeed, in practice, due to numerical approximations, we cannot guarantee that the eigenvalues of the Laplacian are all distinct and, possible symmetries in the shapes may cause some of them to have multiplicity grater than one. As shown by Mateus et al. [42], when dealing with shape matching, an elegant way to overcome this problem is to use the Laplacian eigenmaps scheme [4], which can be seen as a reduced GPS, and perform a-posteriori alignment of the resulting point cloud embeddings. Matching the eigenfunctions histograms, i.e. their signatures, is the only reliable method for such embeddings alignment [42, 60]. Unfortunately, when the eigenvalues separation is small due to symmetries in the geometries, this method becomes unstable and sensitive to noise. To overcome this issue, in this work we propose the Coupled Laplacian operator to produce a-priori aligned eigenmaps for several registered similar shapes, and perform locally-aware shape matching.
More recently, deep-learning based point-cloud descriptors [25, 72, 6, 77] arose as an alternative to generate local features for 3D sufraces. Here we can distinguish three main categories depending on the backbone architecture they employ. Convolutional Neural Networks (CNNs) are often used on point clouds projected into 2D depth images [17, 30] or directly on 3D voxels [25, 77, 30]. For example, 3DMatch [77] aims to learn a local volumetric patch descriptor for establishing correspondences between partial 3D data. Secondly, to work directly on raw point cloud data, PointNet and PointNet++ have been proposed aiming to learn rotation and permutation invariant features [49, 50]. Based on PointNet, several learned descriptors have been proposed [15, 14, 75, 76]. Among them, PPFNet [15] and PPF-FoldNet [14] try to improve the feature representation of PointNet by incorporating global context and point-pair features. Nevertheless, the lack of convolutional layers in these models limits the learning of local geometries. Finally, transformers-based descriptors, such as Deep Closest Point [68], have been recently proposed trying to exploit the attention mechanisms to capture shapes and surfaces intrinsic characteristics. Arguably, training a deep-learning model, especially if based on transformers, it is not always feasible in terms of training samples required to achieve good performances. Furthermore, in supervised algorithms [77, 25, 15, 14], paired 3D patches, such as in the 3DMatch dataset [77], are needed to train the models. For this reason, the 3D rigid point cloud matching method we propose here does not need training and utilizes only the similarity among properly aligned spectral embeddings to relate an unseen target shape to the given references.
3 Method
Figure 2: Proposed workflow of the Coupled Laplacian applied to proximal femur shapes. The $N$ sources are registered to the target using a rigid or affine registration. After that, cross-connection are added between each target-source pair (for simplicity, in the zoom the shapes are not overlapping) and the Coupled Laplacian is computed on the global graph. Its eigendecomposition leads to aligned spectral embeddings of the input geometries that can be used for shape matching.
Graph Laplacian. A 3D point-cloud $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ can be treated as a connected undirected weighted graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ is the nodes set and =$\mathcal{E}=\{\boldsymbol{e}_{ij}\}$ is the edge set. The latter is generally obtained though the construction of a nearest neighbors graph [4]. To that purpose, in this work, we consider the $k$-Nearest Neighbour ($k$-NN) approach, that is, node $j$ is connected to node $i$ if $\boldsymbol{x}_{i}$ is among the $k$ nearest neighbours of $\boldsymbol{x}_{j}$. Hence, we can build a weighted adjacency matrix, $\boldsymbol{W}=\{w_{ij}\}$, which stores the connections between nodes. In particular, in spectral graph theory, a Radial Basis Function (RBF) is commonly used as weight function for the edges between $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$, as
$w_{ij}=\exp\left(-\frac{\text{d}^{2}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}{\sigma^{2}}\right)$ (1)
[…]
$\textbf{L}\boldsymbol{\phi}_{i}=\lambda_{i}\boldsymbol{B}\boldsymbol{\phi}_{i},\;\;i=0,\cdots,n.$ (2)
where $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is generally set as $\boldsymbol{D}$, and $\lambda_{i}\leq\lambda_{i+1}$. Finally, eigenmaps are simply eigenvectors sub-spaces, generated by leaving out $\boldsymbol{\phi}_{0}$, corresponding to $\lambda_{0}=0$ and using the next $m$ eigenvectors for embedding graph nodes in an $m$-dimensional space, $\boldsymbol{x}_{i}\rightarrow[\boldsymbol{\phi}_{1}(i),\cdots,\boldsymbol{\phi}_{m}(i)]$
Coupled Laplacian. The complete pipeline to generate aligned Laplacian embeddings through graph coupling is illustrated in Fig. 2. Let $\mathcal{G}^{\mathcal{T}}$ be the graph corresponding to a target 3D point cloud and $\mathcal{G}^{S_{1}},\cdots,\mathcal{G}^{S_{N}}$ the ones obtained from $N$ different sources. We construct a global graph, or coupled graph, $\mathcal{G}^{C}$, by adding cross-connections separately between the vertices of the source, $\mathcal{V}^{\mathcal{T}}=\{\boldsymbol{x}^{\mathcal{T}}_{i}\}_{i=1}^{n^{\mathcal{T}}}$, and each of the reference shapes, $\mathcal{V}^{S_{k}}=\{\boldsymbol{x}^{S_{k}}_{i}\}_{i=1}^{n^{S_{k}}}$ with $k=1,\cdots,N$. To include meaningful cross-connections, we first perform a rigid [20, 21, 45] or affine [45] registration of the source geometries to better match the target. The latters are preferred to a non-rigid registration because, when dealing with the identification of fine local variations, they do not change the relative position of the points inside a point-set and, therefore, the deformations are kept almost unchanged. After that, a sub-set of vertices, $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ , is stochastically extracted from the target geometry and its nearest correspondence is searched in each of the aligned reference point clouds, such as
$F^{S_{k}}=\{\boldsymbol{f}^{S_{k}}_{j}:\min_{\boldsymbol{x}^{S_{k}}_{i}\in\mathcal{V}^{S_{k}}}\text{d}^{2}(\boldsymbol{x}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{j})|\boldsymbol{f}^{\mathcal{T}}_{j}\in F^{\mathcal{T}}\}$ (3)
with $k=1,\cdots,N$. In this way, there is no constraint on the original number points comprising each shape nor on their initial coordinates system, e.g. camera coordinates. Furthermore, the cardinality of the target sub-set, is chosen as a fraction of its total points, $|F^{\mathcal{T}}|=l\cdot n^{\mathcal{T}}$, where $0<l\leq 1$. The cross-connection are therefore added for each pair $(\boldsymbol{f}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{i})$, $1\leq i\leq|F^{\mathcal{T}}|$, between the target and the $k$-th source. Finally, we arrange the vertex indices of the coupled graph such that they are grouped for each individual geometry, and the whole weighted adjacency matrix can be computed as in Eq. 1.
We define the Laplacian matrix of a global graph, constructed as describe above, as the Coupled Laplacian, $\boldsymbol{L}^{C}\in\mathbb{R}^{n\times n}$, where $n=n^{\mathcal{T}}+\sum_{k=1}^{N}n^{S_{k}}$. Thanks to the vertex ordering of the coupled graph, the solution of the generalized eigenproblem of Eq. 2 applied to the Coupled Laplacian with $\boldsymbol{B}^{C}=\text{diag}([\boldsymbol{D}^{\mathcal{T}},\boldsymbol{D}^{S_{1}},\cdots,\boldsymbol{D}^{S_{N}}])$, which is the stack of uncoupled diagonal degree matrices, gives coupled eigenvectors $\{\boldsymbol{\phi}^{C}_{i}\}_{i=0}^{n}$. These eigenvectors can then be split into each single component of the coupled graph as
$\boldsymbol{\phi}^{C}_{i}=[\boldsymbol{\phi}^{\mathcal{T}}_{i},\;\boldsymbol{\phi}^{S_{1}}_{i},\;\cdots,\;\boldsymbol{\phi}^{S_{N}}_{i}]^{T}.$ (4)
The eigenspaces restricted to the single geometries, obtained thought the Coupled Laplacian, are intrinsically aligned up to a certain component, depending of the factor $l$, and the eigenvalues ordering issue is automatically solved since the split eigenvectors are associated to the same eigenvalues. The proof of this property for the ideal case of perfect match and more theoretical aspects of the Coupled Laplacian are reported in the Supplementary Material.
Shape Matching with Eigenmaps. The Coupled Laplacian allows the generation of $m$-dimensional embeddings aligned for the graph vertices of the target and reference shapes, $\boldsymbol{\Phi}^{\mathcal{T}}=(\boldsymbol{\phi}^{\mathcal{T}}_{1},\cdots,\boldsymbol{\phi}^{\mathcal{T}}_{m})\in\mathbb{R}^{n_{T}\times m}$ and $\boldsymbol{\Phi}^{S_{k}}=(\boldsymbol{\phi}^{S_{k}}_{1},\cdots,\boldsymbol{\phi}^{S_{k}}_{m})\in\mathbb{R}^{n_{S_{k}}\times m}$ for $1\leq k\leq N$, respectively. Therefore, a comparison of multiple geometries in this higher dimensional spectral space yields to a proper consideration of local structures. We consider in the following two possible applications of the coupled eigenmaps: (1) a global shape matching score and (2) local similarity scores. The first one can be obtained by measuring the similarity between the aligned eigenspaces through the Grassmann distance, $d_{G}(\cdot,\cdot)$ [74]. Nevertheless, it cannot be computed directly, since we consider point-sets of arbitrary size and we do not have point-to-point correspondences. Hence, we restrict the distance computation only to the set of cross-connected vertices, $F^{\mathcal{T}}$ for the target and $F^{S_{k}}$ for the $k$-th reference. In order to get the reduced basis restricted only to those coupled points, we perform a QR decomposition of the restriction of the eigenmodes to the cross-points. The best matching source is therefore given by
$\operatorname*{arg\,min}_{k}d_{G}(\boldsymbol{Q}^{\mathcal{T}},\boldsymbol{Q}^{S_{k}}).$ (5)
where $\boldsymbol{Q}^{\mathcal{T}}$ and $\boldsymbol{Q}^{S_{k}}$ are obtained from the QR factorization of $\boldsymbol{\Phi}^{\mathcal{T}}(F^{\mathcal{T}},:)$ and $\boldsymbol{\Phi}^{S_{k}}(F^{S_{k}},:)$, respectively. This first approach is used in the following to solve the BSE task by using two source shapes, i.e. the reference bone of known side and its mirrored version.
On the other hand, to obtain point-wise similarity scores we propose to compare the $m$-dimensional embeddings of the cross-connection vertices by using the cosine distance function. In this case, we interpret the distance value as the probability of local structural difference, where 0 means that the two compared points have the same local structure.
4 Experiments
4.1 Experiment settings
Table 1: Summary of the point cloud bone structures collected for the BSE benchmark. L and R stands for Left and Right, respectively, and S for Sheep.
Dataset Femur Hip Tibia Fibula
L
R
L
R
L
R
L
R
Human
Fisher et al. [19]
18
19
20
20
SSM-Tibia [32]
30
30
ICL [47]
35
35
35
35
35
35
S
Internal [40]
18
18
18
18
Total
71
72
20
20
53
83
35
65
Table 2: Anomaly localization results. The area under the PRO curve is reported for an integration limit of $0.3$ for each evaluated method and dataset category. GAN, AE and VM results are provided by Bergmann et al. [7], and, for a fair comparison, we include the results obtained by restricted GPS [55], eigenfunctions (Hist) [42, 60] and Euclidean matching. The subsctipt on a method indicates the number of used eigenmaps. The overall best performing methods are highlighted in boldface, while the bests for each category are underlined.
Method
Bagel
Cable Gland
Carrot
Cookie
Dowel
Foam
Peach
Potato
Rope
Tire
Mean
3D RGB
VM
0.388
0.321
0.194
0.570
0.408
0.282
0.244
0.349
0.268
0.331
0.335
GAN
0.421
0.422
0.778
0.696
0.494
0.252
0.285 0.362
0.402
0.631
0.474
AE 0.432
0.158
0.808
0.491
0.841 0.406
0.262
0.216
0.716
0.478
0.481
3D Only
GAN
0.111
0.072
0.212
0.174
0.160
0.128
0.003
0.042
0.446
0.075
0.143
AE
0.147
0.069
0.293
0.217
0.207
0.181
0.164
0.066
0.545
0.142
0.203
VM
0.280
0.374
0.243
0.526
0.485
0.314
0.199
0.388
0.543
0.385
0.374
[…]
0.452
0.616
0.695
0.364
0.738
0.471
0.659
0.844
0.647
0.651
0.613
[…]
0.465
0.621
0.690
0.363
0.739
0.480
0.672
0.833
0.653
0.670
0.619
[…]
0.476
0.629
0.703
0.365
0.744
0.473
0.661
0.840
0.647
0.693
0.623
[…]
0.491
0.629
0.698
0.351
0.746
0.501
0.663
0.841
0.652
0.695
0.627
Euclidean
0.655
0.631
0.743
0.615
0.803
0.528
0.726
0.875
0.762
0.695
0.703
$\textbf{Ours}_{100}$
0.669
0.642 0.808
0.714
0.812
0.582
0.748
0.897
0.750
0.733
0.736
$\textbf{Ours}_{200}$ 0.702
0.630
0.728
0.735 0.812 0.701 0.780 0.914 0.767
0.713
0.748
Table 3: Accuracy of human BSE for each bone structures. All the matching methods are applied after RANSAC registration. The overall best performing methods are highlighted in boldface.
Method
Femur
Hip
Fibula
Tibia
Mean
Hausdorff
73.52
97.95
59.30
72.52
75.82
Chamfer
71.64
98.65
64.62
74.47
77.35
FPFH [57]
68.88
96.67
66.69
78.35
77.65
$\textbf{Ours}_{20}$
78.79
98.78
71.28
78.46
81.83
$\textbf{Ours}_{10}$ 79.68
97.76
73.47 78.66 82.39
Bone Side Estimation. The proposed global shape matching score is used to perform automatic BSE, that is, given a reference bone of known side we want to assess the side of another bone, referred as target. Assuming general unknown initial frames, we used the Principal Component Analysis (PCA) to generate a mirrored version of the reference surface. Namely, we identified that the bilateral symmetry of human and animal bodies is equivalent to a mirroring around the second principal component on the vast majority of bones, such as the ones considered in this study. Hence, given the three point clouds, target, source and mirrored source, the Coupled Laplacian of Fig. 2 can be applied. More specifically, we selected the Random Sample Consensus (RANSAC) [20], preceded by a spectral scaling, as rigid registration method. The scaling is performed using only the information carried by the Fiedeler vector, i.e. the eigenvector $\boldsymbol{\phi}_{1}$ corresponding to the first non-zero eigenvalue $\lambda_{1}$. A rough estimation of the bones length, independently of the Euclidean frame, is given by the distance between the points corresponding to minimum and maximum values of the Fiedler vectors. Thanks to that, one of the two bones can be scaled in order to match the length of the other and improve the RANSAC registration. The three aligned $m$-dimensional embeddings, derived from the eigendecomposition of the Coupled Laplacian, are then used as in Eq. 5 to retrieve the best matching source and, consequently, the target side. Note that, the proposed BSE method is fully independent of the target and reference frames and therefore can be applied directly on the segmentation obtained from a medical image, without any previous knowledge, in a fully automated pipeline. Additionally, in alternative, one can using two different bone references, one left and one right, avoiding the mirroring step, or apply the mirroring on the target shape while letting the source unchanged. Full details of the described algorithm are provided in the Supplementary Material.
We generated a benchmark to test the BSE task by collecting several bone surface point cloud data from public datasets [19, 32, 47]. Namely, from Fisher et al. [19] we extracted Femur and Hip structures, from SSM-Tibia [32] only right Tibia and Fibula, and from the Imperial College London (ICL) [47] we collected Femur, Tibia and Fibula point clouds. The composition of the benchmark is shown in Tab. 1 and the detailed instructions to collect and prepare it for testing are provided in the Supplementary Code (or GitHub). Such a variety of shapes, which are acquired using different methods, makes the task more challenging, including some intra-clinic variability. Some additional internal data [40] with knee structures of sheep are used in the Supplementary Material to test the human to animal generalization, and vice-versa, of the proposed BSE technique. In our experiments, for a given bone class, we selected each shape once as source and the others as targets on which the side has to be inferred. Therefore, the average accuracy of all the experiments takes into account the variability of the source shape, and the method robustness can be analysed.
Figure 3: Average BSE accuracy with respect to the fraction of cross-edges used to build the Coupled Laplacian and the number of eigenmaps used for matching.
Figure 4: Graphical comparison of 3D anomaly localization using different numbers of aligned eigenmaps, $m$. Tuning the dimension of the embeddings computed with the Coupled Laplacian, it is possible to decide the extent and size of local surface differences that are detected.
Anomaly Localization. We tested the proposed approach for the 3D anomaly detection task in the MVTec 3D-AD dataset [7]. This recent dataset was designed for the unsupervised detection of anomalies in point clouds of industrially manufactured products. It contains over 4000 high-resolution 3D scans of 10 object categories. In our experiments, we chose one anomaly-free training sample for each class as source and we used the point-wise distances of the aligned eigenmaps between this shape and each target in the test set as anomaly scores. In this case, we performed the registration using affine Coherent Point Drift (CPD) [45], which is more suited for objects like the ones included in the MVTec 3D-AD, e.g. the length and thickness variation among potatoes and carrots samples is better captured by an affine registration rather than rigid. In order to speed-up the computation, we pre-processed the point clouds by removing their flat background with a threshold on the $z$-axis after a 3-dimensional PCA and, without losing generality, we sub-sampled the number of foreground points to a maximum of $13000$. Finally, to compare the obtained anomaly scores with the Ground Truth (GT) image, we projected back the points to the original 2D plane and, if any sub-sampling was performed in pre-processing, we applied a dilation with a structuring element of size the inverse of the sampling factor.
4.2 Experiment results
We show here comparative quantitative and qualitative results of the proposed method, while a study of the robustness with respect to the source shape and more graphical comparisons are provided in the Supplementary Material.
Bone Side Estimation. In Fig. 3, the average accuracy of BSE is depicted based on the fraction of artificial cross-edges added to compute the Coupled Laplacian and the dimension of the spectral embeddings, $l$ and $m$, respectively. We can observe that the more coupled eigenmaps we aim to utilize for computing a global similarity score, the greater the need for additional cross-connections to ensure the reliability of the results. This occurs because, when graphs are weakly coupled with only a few cross-connections, the Coupled Laplacian captures limited intra-shape characteristics and predominantly emphasizes individual local geometries. In fact, when $l=0$ and therefore the coupled graph is not fully connected, the eigenmaps obtained using the Coupled Laplacian are equivalent to the ones computed independently on each single graph, if the eigenvalues are the same (see Supplmentary Material). On the other hand, not too many cross-edges, nor eigenmaps, are needed to obtain meaningful global matching scores and achieve good performances in the BSE task. In Tab. 3 a quantitative comparison of human BSE with respect to other distance measures is reported. For a fair compariosn, all the other metrics are computed after RANSAC registration of source and mirrored source. Hausdorff and Chamfer discrepancies are calculated on the Euclidean coordinates of the points, while FPFH [57] are 33-dimensional vectors describing each point based on histograms of the surface normals. Our method, using $10$ and $20$ eigenmodes with $l=0.5$ fraction of cross-connections, outperforms the other techniques achieving better accuracy both on the single bones and in average. This is because the description provided by the aligned eigenmaps, obtained through Coupled Laplacian, is more aware of local details than the other distance metrics.
Anomaly Localization. Tab. 2 lists quantitative results for each evaluated method for the localization of anomalies. For each category, the normalized area under the Per-Region Overlap (PRO) curve [7] with an upper integration limit of $0.3$, as well as, the mean performance, are reported. Performance of Generative Adversarial Network (GAN), Autoencoder (AE) and Variation Model (VM) on the same test set are provided by the dataset authors [7]. Furthermore, we tested restricted GPS [55], eigenmaps histogram matching [42, 60] and our method, all with the same pre-processing and source shapes, using $100$ and $200$ eigenmaps. In order to obtain denser and more precise anomaly localization, the Coupled Laplacian is built using the whole set of points, after pre-processing, as cross-connections, i.e. $l=1$. Furthermore, we include the result obtained using as anomaly score for each target point the normalized Euclidean distance of the nearest point in the source geometry after registration. Our method, using only 3D information, outperforms all the other techniques in average and on most of the individual objects. Moreover, we obtained better results than deep-learning methods having RGB and Depth in combination as input. Interestingly, the point-wise similarity computed on eigenmaps not properly aligned (GPS and Hist) is worst than just considering Euclidean distances between points of registered shapes, making worthless the computation of spectral embeddings. In Fig. 4 we compare the qualitative results obtained using different numbers of eigenmaps to score the point-wise similarities. Using a smaller $m$, the proposed technique is prone to individuate only small regions with highly dissimilar local geometries. By increasing the size of the spectral embeddings, we can identify larger and more subtle surface differences, even if they are not necessarily classified as anomalies in the GT. For instance, the potato and bagel surfaces have some natural irregularities, with respect to the source, that are not highlighted using $100$ and $200$-dimensional embeddings, but are instead detected with $300$ eigenmaps. This concept is linked to the modal length, which we define in the Supplementary Material. The trade-off between number of maps used and extent of surface differences detected is interesting to tune the proposed method to other tasks requiring specific attention to identify surface dissimilarities.
Figure 5: Comparison of the distance between source and target, in the 3D anomaly detection task, using the Coupled Laplacian technique and Euclidean distance after RANSAC rigid alignment.
Ablation Study. To further motivate the use of the Laplacian spectral space over the Euclidean space, in Fig. 5 we show some comparison of the distance between source and target, after RANSAC rigid alignment, in the two spaces, the latter obtained using the proposed Coupled Laplacian operator with $m=200$. In the first and second row the better resistance to noise and outliers in the spectral space is highlighted. The bending of the cable gland, i.e. the anomaly, is not detected in the Euclidean space because of an additional segment of points on the back of the target having the largest distance to the source. On the other hand, with the aligned eigenmaps comparison, the anomaly, as well as the noise, is correctly localized. This is because both are local differences, with respect to reference, and, therefore, they can be easily captured in the Laplacian space. In addition, in the third row, we show that our technique provides a sharper anomaly split of the cookie point cloud, which simplifies the anomaly detection task. Nevertheless, the usage of the proposed landmarks-free Coupled Laplacian is limited to tasks in which a correspondence between target and source shapes can be defined by mean of a registration process. Problems including graph isomorphism, such as articulated shape matching [42, 60], or non-rigid shape matching, can be still solved using our method, without relying on the rigid alignment, if some cross-connection are defined a-priori. In this case, landmarks and their neighbouring points would be most likely the easiest option to connect separate graphs and obtain aligned spectral embeddings for further processing. Moreover, a wrong registration of the shapes, due to randomness and errors in the selected registration algorithm, may affect the alignment of the $m$-dimensional embeddings derived from the Coupled Laplacian. Even if the probability of such event for similar geometries is relatively low and the adjacency matrix is already weighted with RBF, a double-check can be added in the pipeline to verify the correct registration [13].
5 Conclusions
We presented a versatile method to perform 3D rigid point cloud matching, both globally and locally, by using aligned eigenspaces for two or more similar shapes without prior knowledge on the local frames or specif markers on them. The introduction of the Coupled Laplacian operator enables the generation of aligned eigenmaps relying on cross-connections added between graphs of registered geometries. We introduced a novel task consisting in the automatic detection of a bone side, i.e. Bone Side Estimation (BSE), and we proposed a benchmark to test it. Leveraging global similarities of eigenmaps derived from the Coupled Laplacian, we define a full pipeline to perform BSE on arbitrary bone surfaces segmented from a generic medical image. Moreover, we tested the ability of the proposed method in capturing surface differences by performing local shape matching for the 3D anomaly detection task on the MVTec 3D-AD dataset. The proposed technique outperforms other methods on the two tasks, therefore capturing better both global and local similarities, thanks to the matching performed on correctly aligned spectral spaces. Beside the applications showcased here, we believe that our work can help several tasks in many fields in which the localization of local difference is crucial for matching, detection or retrieval. Future extensions could explore generalizations of the technique, not only for rigid matching but also to address challenges such as graph isomorphism. This approach opens new avenues for advancing the State-of-The-Art in 3D shape analysis and matching techniques.
6 Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 945304-Cofund AI4theSciences hosted by PSL University. This work was granted access to the HPC/AI resources of IDRIS under the allocation 2022-AD011013902 made by GENCI.
Supplementary Material
7 Coupled Laplacian
We present here some theoretical aspects of the Coupled Laplacian, introduced in Sec. 3, which we proposed to generate aligned spectral embeddings for multiple registered point clouds.
Zero Cross-Connections. The coupled Laplacian matrix has the following structure:
$\boldsymbol{L}^{C}=\boldsymbol{L}^{U}+\boldsymbol{L}^{+}$ (6)
where $\boldsymbol{L}^{U}$ is the Laplacian matrix of the global graph without any cross-connection between $\mathcal{G}^{\mathcal{T}}$ and $\mathcal{G}^{S_{k}}$, for $k=1,\cdots,N$. Hence, $\boldsymbol{L}^{+}$ is the Laplacian matrix of the global graph restricted to its cross-connections between sub-graphs. Moreover, $\boldsymbol{L}^{+}$ enforces the matching constraint in the global eigenproblem.
By construction of the matrix $\boldsymbol{B}^{C}$ (see Sec. 3), eigenvectors related to $(\mathcal{G}^{S_{k}})_{k=1,\cdots N}$ and $\mathcal{G}^{\mathcal{T}}$, that have the same eigenvalue $\lambda$, can be stacked into a global eigenvector $\boldsymbol{\phi}^{U}$ such that the following property holds
$\boldsymbol{L}^{U}\>\boldsymbol{\phi}^{U}=\lambda\>\boldsymbol{B}^{C}\>\boldsymbol{\phi}^{U}$ (7)
where the above system is block diagonal. Therefore, solving the Coupled Laplacian eigenproblem without cross-connections, i.e. $l=0$ and then $\boldsymbol{L}^{+}=\boldsymbol{0}$, it is equivalent of solving separate eigenprobems, when the eigenvalues are the same, for each single component of the global graph. Otherwise, if eigenvalues are not the same, the solution of the coupled eigenproblems with $\boldsymbol{L}^{+}=\boldsymbol{0}$ will likely give separate eigenvectors for each sub-graph, that is, vectors with non-zero values only for indices corresponding to a single component.
Ideal Matching. In the ideal case of perfect matching, when source graphs bare copies of the target graph, whether or not vertices are randomly reordered, the set of $n^{\mathcal{T}}$ coupled eigenvectors contains $N+1$ copies of target’s eigenvectors. For the sake of simplicity, in the following proof, the vertices in the copies of the target graph are not reordered. Therefore, $\mathcal{G}^{S_{k}}=\mathcal{G}^{\mathcal{T}}$ and $F^{S_{k}}=F^{\mathcal{T}}$, for $k=1,\cdots,N$. Let $\boldsymbol{\eta}^{C}_{i}\in\mathbb{R}^{(N+1)n^{\mathcal{T}}}$ be the vector in which $N+1$ copies of an eigenvector $\boldsymbol{\phi}^{\mathcal{T}}_{i}$ are stacked. Then, from Eq. 6 and Eq. 7 we have,
$\boldsymbol{L}^{C}\>\boldsymbol{\eta}^{C}_{i}-\lambda^{\mathcal{T}}_{i}\>\boldsymbol{B}^{C}\>\boldsymbol{\eta}^{C}_{i}=\boldsymbol{L}^{+}\>\boldsymbol{\eta}^{C}_{i}$ (8)
where non-zero rows in the right hand side term are related to cross-connections between a vertex of $\mathcal{G}^{\mathcal{T}}$ and its copies on the references, at the same location. Hence, the weight function is equal to one, for each of the cross-connections, and
$\forall j\in F^{\mathcal{T}}\quad(\boldsymbol{L}^{+}\>\boldsymbol{\eta}^{C}_{i})_{j}=\sum_{k=1}^{N}\phi^{\mathcal{T}}_{ji}-\phi^{S_{k}}_{ji}.$ (9)
Since $\boldsymbol{\phi}^{S_{k}}_{i}=\boldsymbol{\phi}^{\mathcal{T}}_{i}$ for $\mathcal{G}^{S_{k}}=\mathcal{G}^{\mathcal{T}}$, the right hand side term of Eq. 8 is equal to zero and $(\boldsymbol{\eta}^{C}_{i})_{i=1,\cdots,n^{\mathcal{T}}}$ are eigenvectors of the Coupled Laplacian containing $N+1$ copies of a target graph.
In addition to the above property, when target graph has eigenvalues with multiplicity higher than one, the coupled term $\boldsymbol{L}^{+}\>\boldsymbol{\phi}^{C}_{i}=0$ enforces the shape matching of eigenvectors that do not belong to the kernel of the matrix $\boldsymbol{L}^{+}$.
Penalization Term. The computation of matching vectors can be penalized in the initial eigenvectors by incorporating a coefficient $\alpha>1$ into the Laplacian matrix for the coupled graph, as follows:
$\boldsymbol{L}^{C\alpha}=\boldsymbol{L}^{U}+\alpha\>\boldsymbol{L}^{+}$ (10)
The larger $\alpha$, the higher the eigenvalues of non matching eigenvectors such that $\|\boldsymbol{L}^{+}\>\boldsymbol{\phi}^{C}_{i}\|>0$.
Modal Length. Let $\boldsymbol{\phi}_{k}$ be the $k$-th eigenvector of a single graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, its gradient on the graph can be expressed as
$||\nabla\boldsymbol{\phi}_{k}||^{2}_{\mathcal{G}}=\sum_{(i,j)\in\mathcal{E}}\left(\frac{\phi_{ik}-\phi_{jk}}{d(\boldsymbol{x}_{i},\boldsymbol{x}_{j})}\right)^{2}$ (11)
where the sum is over the set of edges $\mathcal{E}$ and $d(\cdot,\cdot)$ is the Euclidean distance. We define the modal length as the ratio between the norm of the eigenvector and its gradient
$L_{k}=\frac{||\boldsymbol{\phi}_{k}||}{||\nabla\boldsymbol{\phi}_{k}||_{\mathcal{G}}}$ (12)
with $L_{0}=+\infty$, because $\boldsymbol{\phi}_{0}$ is constant, and $L_{k}>L_{k+1}$ for $k\neq 0$. Note that the unit of $L_{k}$ is the same as the unit given by the Euclidean distance of points in the graph. Therefore, the modal length can give an insight about the number of eigenmpas, $m$, to produce with the Coupled Laplacian, based on the extent of surface differences we want to detect. Intuitively, representing differences at small scales is easier than at large ones. Therefore, the more modes are used, the larger the range of defect sizes that can be represented.
Data: Source surface and side - $\left(\mathcal{V}^{S},\mathrm{S}^{S}\right)$
         Target surface - $\mathcal{V}^{\mathcal{T}}$
         Neighbours - $k$
         Cross-Fraction - $l$
         Eigenmaps - $m$
Result: Target Side - $\mathrm{S}^{\mathcal{T}}$
$\boldsymbol{W}^{S}\leftarrow\text{Adjacency}(\mathcal{V}^{S},k)$;
$\boldsymbol{W}^{\mathcal{T}}\leftarrow\text{Adjacency}(\mathcal{V}^{\mathcal{T}},k)$;
$L^{S},L^{\mathcal{T}}\leftarrow$ Bone lengths from Fiedler vectors;
$\alpha\leftarrow L^{S}/L^{\mathcal{T}}$ ;
  normal-$\triangleright$ Scale factor
$\mathcal{V}^{\mathcal{T}}\leftarrow\alpha\times\mathcal{V}^{\mathcal{T}}$;
$\mathcal{V}^{S}_{M}\leftarrow$ Mirror Source using PCA;
$\mathcal{V}^{S},\mathcal{V}^{S}_{M}\leftarrow$ RANSAC registrations to $\mathcal{V}^{\mathcal{T}}$;
Select target cross-connections set $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ Find sets $F^{\mathcal{S}}\subset\mathcal{V}^{S}$ and $F^{\mathcal{S}}_{M}\subset\mathcal{V}^{S}_{M}$ as in Eq. 3 $L^{C}\leftarrow\text{Coupled Laplacin}(W^{\mathcal{T}},W^{S},F^{\mathcal{T}},F^{\mathcal{S}},F^{\mathcal{S}}_{M})$;
$\{\lambda_{i}^{C}\}_{i=0}^{m}$, $\{\boldsymbol{\phi}_{i}^{C}\}_{i=0}^{m}\leftarrow$ Eig$(L^{C},m)$;
$\boldsymbol{\Phi}^{\mathcal{T}},\boldsymbol{\Phi}^{S},\boldsymbol{\Phi}^{S}_{M}\leftarrow$ Split($\{\boldsymbol{\phi}_{i}^{C}\}_{i=0}^{m}$);
QR decompose $\boldsymbol{\Phi}^{\mathcal{T}}(F^{\mathcal{T}},:),\boldsymbol{\Phi}^{S}(F^{S},:),\boldsymbol{\Phi}^{S}_{M}(F^{S}_{M},:)$;
if $d_{G}(\boldsymbol{Q}^{\mathcal{T}},\boldsymbol{Q}^{S})\leq d_{G}(\boldsymbol{Q}^{\mathcal{T}},\boldsymbol{Q}^{S}_{M})$ then
      $\mathrm{S}^{\mathcal{T}}$ = $\mathrm{S}^{S}$ ;
        normal-$\triangleright$ Same side
else
      $\mathrm{S}^{\mathcal{T}}$ = $\neg\mathrm{S}^{S}$ ;
        normal-$\triangleright$ Opposite side
end if
Algorithm 1 Spectral-Based Bone Side Estimation (BSE) using aligned embeddings from the Coupled Laplacian.
8 Spectral-Based BSE
In Algorithm 1 we provide full details on the spectral-based algorithm for BSE introduced in Sec. 4.1. It takes as input the set of points, i.e. graph vertices, representing source and target surfaces , $\mathcal{V}^{S}$ and $\mathcal{V}^{\mathcal{T}}$, respectively, and the target side, $\mathrm{S}^{S}$. Additional input parameters are the number of neighbours used to create the $k$-NN graphs from the point clouds, the fraction of target nodes used to create cross-connection, $l$, and the dimension of the embeddings generated through the Coupled Laplacian, $m$.
As first step, the single adjacency matrices are built from the set of vertices using an RBF kernel, as in Eq. 1. After that, the length of each shape is computed only with the information carried by the Fiedler vector, $f$, as follows
$L=||x_{\text{max}}-x_{\text{min}}||^{2}$ (13)
where $x_{\text{max}}=x_{\operatorname*{arg\,max}{f}}$ and $x_{\text{min}}=x_{\operatorname*{arg\,min}{f}}$ are the points corresponding to the maximum and minimum values of the Fiedler vector, respectively. In this way, the input surfaces do not need to lie on the same Euclidean frame in order to have comparable length measurements. Even if Fiedler vectors of different shapes may have opposite sign, the inversion of maximum and minimum does not affect the bone length calculation. Note that, as extension for a more precise length computation, it is also possible to select the $M$ maximum and minimum points of the Fiedler vector and compute $x_{\text{max}}$ and $x_{\text{min}}$ as the barycenters of those points. The ratio between source and target lengths gives a scaling factor, $\alpha$, that we use to scale the target surface in order to match the length of the reference. The mirrored version of the source bone, $\mathcal{V}^{S}_{M}$, is then generated by flipping it around the point cloud second principal component and both versions of the reference are registered to the target with RANSAC algorithm [20]. Interestingly, the adjacency matrix of the two references, before and after registration, is the same as the original one and, therefore, we can avoid its computation multiple times.
Figure 6: Graphical representation of the structure of the coupled adjecency matrix. $\boldsymbol{W}^{S}$ and $\boldsymbol{W}^{\mathcal{T}}$ are the source and target adjacency matrices, while $F^{\mathcal{T}}$, $F^{S}$ and $F^{S}_{M}$ are the set of nodes for cross-connections of target, source and mirrored source. Note that the nodes are sorted in groups such that the eigenmaps derived from the Coupled Laplacian can be easily split into the single components.
Figure 7: Box plot of the accuracy of various methods for BSE depending on the source shape. Each measurement is the accuracy obtained using a different source shape among the ones available in the proposed benchmark. The median is highlighted with the notch while the mean is represented by the dashed line.
After the described pre-processing, the core of the algorithm based on the Coupled Laplacian takes place. The sub-set of target nodes used for cross-connections is stochastically selected and the corresponding points on the two versions of the reference are queried as in Eq. 3. Hence, the three graphs can now be connected and the Coupled Laplacian computed from the whole weighted adjacency matrix. For more clarity, the structure of the coupled graph adjacency matrix is depicted in Fig. 6. Therefore, the first $m$ coupled eigenmaps are computed with a given eigensolver and split into the components of each of the three single shapes, as in Eq. 4. Finally, the Grassman distances, $d_{G}(\cdot,\cdot)$, of the QR normalized aligned eigenmaps restricted to the cross-connection sets are compared in order to predict the target side. The latter is equal to $\mathrm{S}^{S}$ if the Grassman distance between target and reference is lower than the distance between target and mirrored reference, otherwise it will be the opposite body side.
9 Robustness to Source Variation
In this section we briefly discuss the robustness with respect to the selected source shape of the methods tested in Sec. 4.2.
Table 4: Minimum and maximum accuracy of human BSE for each bone structures using different sources. All the matching methods are applied after RANSAC registration. The overall best performing methods are highlighted in boldface.
Method
Femur
Hip
Fibula
Tibia
Best Hausdorff
89.62
100.0
75.76
86.87
Chamfer
85.85
100.0
86.87
90.91
FPFH [57]
83.96
100.0
83.84
93.94
$\textbf{Ours}_{20}$ 93.40 100.0
89.90
94.95
$\textbf{Ours}_{10}$ 93.40 100.0 91.92
93.94
Worst Hausdorff
51.89
89.74
39.39
51.51
Chamfer
54.72
92.31
35.35
38.38
FPFH [57]
34.34
87.18
34.34
41.41
$\textbf{Ours}_{20}$
49.06
94.87
38.38
39.39
$\textbf{Ours}_{10}$ 57.55
84.62
43.43
45.45
Bone Side Estimation. Fig. 7 shows the variation of the BSE accuracy, with respect to the source bone, for different bones and different methods. We recall that in each experiment one bone is chosen as source and the side is inferred on all the other bones of the same category. Therefore, we automatically have $N_{b}$ accuracy measurements for each class, where $N_{b}$ is the number of samples of the bone class $b$. The average accuracy of these experiments is reported in Tab. 3 of the main manuscript. While median and average accuracy are generally higher using our technique in all the bones, the values spread is similar between methods. Additionally, we report in Tab. 4 the maximum and the minimum accuracies obtained using different sources. We can observe that the maximum accuracy is always achieved with our method, both using $10$ or $20$ eigenmaps for matching. Moreover, it is also better in the worst case for all the bones classes except for the tibia, for which the Hausdorff distance seems to be better. Nevertheless, as also shown in Fig. 7, there is no much difference in the span between best and worst cases among different methods. This suggests that, to improve the performance of our BSE technique, and also of other methods, a proper choice of source bone is essential.
Figure 8: Area under the PRO curve (AUC-PRO) with an upper integration limit of 0.3 for each category of the MVTec 3D-AD dataset. Each score, represented by a dot, is computed with a different source shape selected from the train set. The median is highlighted with the notch while the mean is represented by the dashed line.
Figure 9: Graphical comparison of 3D anomaly localization using different source shapes on different objects. The target example with RGB colors and the GT anomaly are shown on the left and right, respectively. In the middle $6$ different results obtained changing the anomaly-free source are reported.
Anomaly Localization. We tested our method on the 3D anomaly localization task using $20$ different anomaly-free sources randomly drawn from the train set of the MVTec 3D-AD dataset. For each train sample, and for each category, we computed the normalized area under the PRO curve with an upper integration limit of 0.3 and their distribution is shown in Fig. 8. All the experiments are performed using the Coupled Laplacian with $m=200$ and $l=1$. We can observe that the results are much more stable and less spread than in the BSE task. This is likely due to the fact that different samples of the same industrially manufactured product are very similar and, therefore, all equally good as source shape. On the other hand, bones may vary in size and slightly in form from patient to patient. Hence, the selection of a good representative source is essential to enhance the capabilities of the proposed matching technique.
Moreover, in Fig. 9 we compare graphically the results obtained on some test samples using different anomaly-free sources. We can observe the strength of the proposed method which always correctly localizes the anomalies, even with different sources. In opposition, some false positives are sometimes present, especially on the borders of the shapes. This weakness is due to fact that the considered point clouds are not closed surfaces, since they are acquired with objects placed on a flat surface. Therefore, the points on the edges between object and surface are likely to contain noise which may produce different coupled embeddings leading to a detection of local surface differences. Interestingly, in the peach sample, some anomalies are located in the area of the steam end because only some anomaly-free sources has it. The peach test example shown in Fig. 9 does not have the steam and therefore, when compared with a source that has it, a surface difference is highlighted in that specific area.
Figure 10: Graphical comparison of 3D anomaly localization using different methods on different objects. The target example with RGB colors and the GT anomaly are shown on the left and right, respectively. In the middle the results obtained with FPFH, Euclidean distance and distance between aligned spectral spaces are reported.
10 Graphical Comparison
In this section we first provide more graphical comparisons of different methods on the anomaly localization task. After that, we show the difference between eigenmodes produced with the Coupled Laplacian with respect to the ones generate independently on the single geometries.
In Fig. 10 the anomaly scores obtained using FPFH features comparison [57], Euclidean distance and Coupled Laplacian, with $m=200$ and $l=1$, on different samples are compared. For a fair comparison, all the techniques are performed with the same anomaly-free source sample and after affine CPD [45] source to target registration. We observed that FPFH tends to overestimate anomalies, whereas the Euclidean distance underestimates them, leading to misleading detections. In contrast, our method achieves a better trade-off and improves the localization of anomalies. With reference to the modal length of Eq. 12, decreasing the number of modes used for spectral comparison we will obtain similar results to the Euclidean method, while increasing it, the localization will tend to the one of FPFH. Once again, the parameter $m$ plays a crucial role in our method, helping us select surface differences that are tailored to the specific task at hand.
Figure 11: Graphical comparison of the eigenmaps obtained through Coupled Laplacian and with independent eigendecomposition of point cloud graphs, i.e. uncoupled eigenmaps. The first $6$ eigenmodes are shown in both cases for source, target and mirrored target in the case of femur bones.
Figure 12: Graphical comparison of the eigenmaps obtained through Coupled Laplacian and with independent eigendecomposition of point cloud graphs, i.e. uncoupled eigenmaps. The first $6$ eigenmodes are shown in both cases for source, target and mirrored target in the case of hip bones.
Figure 13: Graphical comparison of the eigenmaps obtained through Coupled Laplacian and with independent eigendecomposition of point cloud graphs, i.e. uncoupled eigenmaps. Some eigenmodes of increasing number are shown for two samples of the MVTec 3D-AD dataset.
In Fig. 11 and Fig. 12 the first $6$ eigenmodes of a femur and hip BSE problem produced with the Coupled Laplacian are shown and compared with the ones obtained thought independent eigendecomposition of target and source, i.e. uncoupled eigenmaps. In both cases the target side is opposite to the source and therefore the best matching coupled eigenmaps are with the mirrored source. This is because its RANSAC registration to the target shape is more precise and so the added cross-connections. On the contrary, the cross-edges between source and target are weaker leading to less precise coupled modes and larger Grassmann distance, which result in a correct BSE. In the uncoupled eigenmaps we can instead observe both the eigenvalue ordering and sing disambiguity of the eigendecomposition. In the femur example, eigenmode $3$ has flipped sign in source and target, while eigenvalues $5$ and $6$ are inverted. Moreover, we can observe slight local variations of modes in the condyles area of the distal femur (highlithed with red circles). Similar observations can be also derived from the uncoupled modes of the hip bones in Fig. 12.
A comparison of higher modes is instead shown for two samples of the anomaly detection task in Fig. 13. In relation to the modal length defined in Eq. 12, it is evident that lower modes represent larger scales compared to higher modes, enabling them to better capture differences. Moreover, the proposed coupling process leads to aligned eigenmaps for source and target shapes. Uncoupled eigenmodes exhibit completely different patterns, rendering them impractical for matching purposes without a reordering process, which complexity increases as the represented scale decreases. On the other hand, modes derived from the Coupled Laplacian exhibit the exact same pattern on both source and target point clouds, making them feasible for direct comparison. However, it is worth noting that as we increase the mode number, there might be an observable weakening in the coupling of the intensities of the eigenvectors, despite the preservation of the underlying patterns. In the computation of similarity scores using aligned embeddings, the observed coupling weakening with increasing mode numbers may not significantly impact the results. In fact, this characteristic could be leveraged to our advantage when determining the truncation number for the eigenspaces, highlighting the adaptability and effectiveness of our method.
Table 5: Accuracy of cross-species BSE for femur and tibia using different methods. All the matching methods are applied after RANSAC registration. S and H stands for Sheep and Human, respectively, and the overall best performing methods are highlighted in boldface.
Method
Femur
Tibia
Mean
S normal-$\rightarrow$ S Hausdorff
94.37
97.14
90.48
Chamfer
94.29
88.57
94.88
FPFH [57]
97.14
97.14
97.14
$\textbf{Ours}_{20}$ 97.30 97.30 97.30
$\textbf{Ours}_{10}$ 97.30 97.30 97.30
H normal-$\rightarrow$ S Hausdorff
72.22
69.44
70.83
Chamfer
69.44
75.00
72.22
FPFH [57]
77.78
75.00
76.39
$\textbf{Ours}_{20}$
69.44
72.22
70.83
$\textbf{Ours}_{10}$ 83.33 77.78 80.56
S normal-$\rightarrow$ H Hausdorff
57.94
63.00
60.47
Chamfer
62.62
66.00
64.31
FPFH [57] 63.89
63.00
63.45
$\textbf{Ours}_{20}$
63.55
60.00
61.78
$\textbf{Ours}_{10}$
63.55
66.00 64.78
Figure 14: Comparison of human and sheep femur and tibia bones. For both of them, the top and lower parts are shown separately to highlight their differences.
11 Cross-Species BSE
We test here the generalization of the proposed BSE algorithm on an internal dataset of sheep femur and tibia bones and we discuss its cross-species capabilities, i.e. when source and target bones are from different species. The latter poses an interesting challenge in the medical field, as tests are often performed on animals before being expanded to humans. The experiments only on sheep bones are performed as described in Sec. 4.1 for the public human benchmark. While, for the cross-species experiments, each bone of one species is chosen once as source to infer the side of the bones of the same category from the other species. The results of these sets of experiments are reported in Tab. 5. We can observe that, the sheep-to-sheep BSE is much more effective then the human-to-human comparison reported in the main manuscript, achieving almost correct estimation in all the sample. This is due to the fact that sheep femur and tibia have less symmetric shapes than their human counterparts, hence making easier the side estimation. Moreover, all the samples in our internal datasets are acquired with the same settings and therefore there is small sample variability, unlike the human benchmark we proposed to collect.
However, despite our method maintaining its superiority, performance drops are noted for cross-species tasks. A clearer understanding of this phenomenon can be obtained from Fig. 14, where human and sheep femur and tibia bones are graphically compared. While the overall shapes are similar, some major differences can be noticed between species. For instance, the sheep proximal femur is similar to the human distal part. Such variations, and other artifacts, may affect the result of the initial registration, leading to wrong cross-connections and miss-aligned eigenmaps. Furthermore, while bones of different species are globally similar, local differences introduce complexities in the BSE task. The global similarity score derived from the Coupled Laplacian is therefore affected by these differences, leading to a more challenging task of aligning and comparing bones across species. Addressing this challenge will contribute to advancing the field of BSE and refining the understanding of global similarities and localized distinctions in diverse bone structures.

### v2

@@
 In some practical scenarios, emphasizing local differences is crucial for accurately identifying a correct match, thereby enhancing the overall robustness and reliability of the matching process.
-Commonly used shape descriptors have several limitations and often fail to provide meaningful local insights on the paired geometries.
+Commonly used shape descriptors have several limitations and often fail to provide meaningful local insights about the paired geometries.
 In this work, we propose a new technique, based on graph Laplacian eigenmaps, to match point clouds by taking into account fine local structures.
-To deal with the order and sign ambiguity of Laplacian eigenmaps, we introduce a new operator, called Coupled Laplacian, that allows to easily generate aligned eigenspaces for multiple rigidly-registered geometries.
+To deal with the order and sign ambiguity of Laplacian eigenmaps, we introduce a new operator, called Coupled Laplacian [^1] Code: https://github.com/matteo-bastico/CoupLap, that allows to easily generate aligned eigenspaces for multiple registered geometries.
 We show that the similarity between those aligned high-dimensional spaces provides a locally meaningful score to match shapes.
-We initially evaluate the performance of the proposed technique in a point-wise manner, specifically focusing on the task of object anomaly localization using the MVTec 3D-AD dataset.
+We firstly evaluate the performance of the proposed technique in a point-wise manner, focusing on the task of object anomaly localization on the MVTec 3D-AD dataset.
 Additionally, we define a new medical task, called automatic Bone Side Estimation (BSE), which we address through a global similarity score derived from coupled eigenspaces.
@@
 Our matching technique, based on Coupled Laplacian, outperforms other methods by reaching an impressive accuracy on both tasks.
-The code to reproduce our experiments is publicly available at https://github.com/matteo-bastico/CoupledLaplacian and in the Supplementary Code.
+footnotetext: Corresponding author: matteo.bastico@minesparis.psl.eu
 1 Introduction
@@
 On the other hand, local descriptors [38, 57] aim to tackle this issue but they generally are sensitive to noise, based on landmarks and they might not capture semantic information [63].
-More recently, deep-learned shape descriptors [72, 6] and deep neural networks for shape matching, based on auto-encoders [73] or transformers [59, 68], have been proposed.
+More recently, deep-learned shape descriptors [72, 6] and neural networks for shape matching, based on auto-encoders [73] or transformers [59, 68], have been proposed.
 Despite their good performances, these methods require a huge amount of annotated data for training, which are hard to collect in fields such as medical imaging [37].
@@
 In this study, we introduce a novel method for rigid 3D point cloud matching, based on spectral Laplacian eigenmaps [4, 23], which focus on local details.
-This technique overcomes the previously mentioned limitations and constraints associated to global and local descriptors and, although it can be seen as a linear graph neural network, it does not require training.
+This technique is designed to overcome the limitations of shape descriptors and, although it can be seen as a linear graph neural network, it does not require training.
 One of the main reasons that combinatorial and geometric Laplacians are often considered for spectral graph processing is that their eigenvectors possess properties like the classical Fourier basis functions [79].
-We exploit these properties to perform locally-aware shape comparison of point clouds, equipped with k-nearest neighbor graph, without relying on additional descriptors.
+We leverage these characteristics to perform locally-aware shape comparison of point clouds, equipped with k-nearest neighbor graph, without relying on additional descriptors.
 Nevertheless, eigenspaces alignment, including both eigenvalue ordering and sign disambiguity of the eigendecomposition, is required to correctly match shapes in the spectral space.
@@
 matching the histograms of the eigenvectors.
-The latter is not robust and frequently fails for highly symmetric geometries, such as bones, with small eigenvalue separation.
+The latter is not robust and frequently fails in possible scenarios with highly symmetric geometries, such as bones, with small eigenvalue separation.
 In this work, we introduce a novel operator, called Coupled Laplacian, designed to simultaneously produce aligned eigenspaces for multiple registered geometries.
@@
 While local similarity is exploited for accurate 3D anomaly detection.
-Hence, the identification of local differences between two shapes, using the proposed matching technique, allows us to perform 3D object anomaly detection in the spectral domain with high accuracy.
-We tested the performances of our method on the MVTec 3D-AD dataset [7], recently proposed for unsupervised point cloud industrial anomaly detection and localization.
-Furthermore, in the context of medical imaging, correctly identifying the side of a bone (left or right side) is crucial for diagnosis, treatment planning, or skeletal analysis [12].
-We refer to the task of determining the side of a Bone or anatomical structure as laterality or Side Estimation (BSE).
+One natural application of the proposed technique, is 3D object anomaly detection thought the identification of local differences between two shapes.
+Hence, we tested our method on the MVTec 3D-AD dataset [7], recently proposed for unsupervised point cloud anomaly detection and localization.
+Furthermore, in the context of medical imaging, correctly identifying the side of a bone (left or right) is crucial for diagnosis, treatment planning, or skeletal analysis [12].
+We refer to this task as Bone Side Estimation (BSE).
 Recent studies on bone landmark detection [18] and statistical shape modeling [31] have highlighted the limitations of manual side identification for ensuring proper functioning.
 To the best of our knowledge, manual bone markings [2] are currently the only technique used for this purpose.
-Consequently, automatic laterality determination arises as an interesting, yet challenging, task to assist the development of fully automated pipelines for medical image analysis, such as patient-specific preoperative planning for Total Knee Arthroplasty (TKA) [34] or Anterior Cruciate Ligament Reconstruction (ACLR) [18, 44].
-Thanks to the bilateral symmetry of the animal and human body [28, 67], i.e.
-right and left sides are mirror shapes of one another, the BSE can be defined as a non-trivial shape matching problem.
-Given a reference, or source, of known side, to assess the side of a target bone, it is enough to find the best match between the source-target pair and its mirrored counterpart, or vice versa.
-The complexity of this task is a consequence of subtle local differences present in mirrored bones.
-This makes it an ideal application of the proposed shape matching technique that utilizes Coupled Laplacian eigenmaps as local geometry descriptors.
-We generate a benchmark for human BSE by extracting surface point clouds of different bones, e.g.
+Consequently, automatic BSE arises as an interesting, yet challenging, task to assist the development of fully automated pipelines for medical image analysis, such as patient-specific preoperative planning for Total Knee Arthroplasty (TKA) [33] or Anterior Cruciate Ligament Reconstruction (ACLR) [17, 44].
+Due to the bilateral symmetry of the animal and human body [27, 67], i.e.
+right and left sides are mirror shapes of one another, the BSE can be defined as a non-trivial chiral shape matching problem.
+Then, given a known reference or source bone, assessing the side of a target bone involves finding the best match between the source-target pair and its mirrored counterpart.
+The complexity of this task arises from subtle local differences in mirrored bones, which makes it a suitable application of the proposed shape matching technique utilizing coupled Laplacian eigenmaps as local geometry descriptors.
+We propose a benchmark for human BSE by extracting surface point clouds of different bones, e.g.
 femur, hip and tibia, from public datasets [18, 31, 47].
-We also discuss, in the Supplementary Material, the cross-species generalization of the proposed matching algorithm from human to animal, and vice-versa, on an internal dataset [40].
+We also discuss, in the Supplementary Material, a non-rigid correspondence application of the proposed matching algorithm, the cross-species BSE (human to animal), on an internal dataset [40].
 Our contributions are summarized as follows:
@@
 2 Related Works
-Early efforts in developing hand-crafted 3D local features for point cloud matching, registration, recognition and retrieval have typically drawn inspiration from 2D descriptors.
+Early efforts in developing hand-crafted 3D local features for point cloud matching, registration and retrieval have typically drawn inspiration from 2D descriptors.
 Several methods, such as Signature of Histograms of OrienTations (SHOT) [58, 66], Rotational Projection Statistics (RoSP) [25] and Unique Shape Context (USC) [66] rely on the estimation of an unique Local Reference Frame (LRF).
@@
 In contrast, LRF-free approaches [10, 56, 57] try to rely just on features that are intrinsically invariant.
-The most common LRF-free features are Fast Point Feature Hisograms (FPFH) [57] which, similarly to SHOT, calculates histograms of surface normals considering their relationships in a local region around a keypoint.
+The most common LRF-free features are Fast Point Feature Hisograms (FPFH) [57] which, similarly to SHOT, calculates histograms of surface normals considering their relationships in a local region around a keypoint and generating $33$-dimensional descriptors.
 Despite the progress made with hand-crafted 3D local features, they encounter difficulties when it comes to deal with issues like point cloud resolutions, noisy data, and occlusions [26].
-On the contrary, spectral-based point cloud descriptors [1, 29, 52, 70, 62, 55, 69] are a category of feature extraction methods that leverage spectral analysis techniques from graph theory and linear algebra to capture the underlying structure and intrinsic geometric properties of point clouds.
-Among them, shape-DNA [52] is a surface and solid descriptor based on eigenvalue analysis of the Laplace-Beltrami operator.
+On the contrary, spectral-based point cloud descriptors [1, 28, 52, 70, 62, 55, 69] are a category of feature extraction methods that leverage spectral analysis techniques from graph theory to capture the underlying structure and intrinsic geometric properties of point clouds.
+Among them, shape-DNA [52] is a surface descriptor based on eigenvalue analysis of the Laplace-Beltrami operator.
 They propose to use the sequence of eigenvalues (spectrum) of the Laplace operator as a fingerprint characterizing the intrinsic geometry of 3D shapes represented as point clouds.
@@
 [70] also proposed to extract a features set consisting of $8$ eigenvalues-based indices for each 3D point of a cloud.
-Furthermore, Heat Kernel Signature (HKS) [62] and Wave Kernel Signatures (WKS) [1] are spectral-based descriptors measuring how heat and wave propagate across a shape, having the eigendecomposition as leading element of the computation.
+Furthermore, Heat Kernel Signature (HKS) [62] and Wave Kernel Signatures (WKS) [1] are descriptors measuring how heat and wave propagate across a shape, having the eigendecomposition as leading element of the computation.
 Scaled eigenvectors evaluated at each point are instead directly exploited by the Global Point Signature (GPS) [55] to represent a point cloud as a set of infinite-dimensional vectors, characterizing each point within the global context of the surface it belongs.
@@
 their signatures, is the only reliable method for such embeddings alignment [42, 60].
-Unfortunately, when the eigenvalues separation is small due to symmetries in the geometries, this method becomes unstable and sensitive to noise.
-To overcome this issue, in this work we propose the Coupled Laplacian operator to produce a-priori aligned eigenmaps for several registered similar shapes, and perform locally-aware shape matching.
-More recently, deep-learning based point-cloud descriptors [25, 72, 6, 77] arose as an alternative to generate local features for 3D sufraces.
-Here we can distinguish three main categories depending on the backbone architecture they employ.
+Recently, Ma et.
+al.
+[39] proposed a canonization algorithm for sign and basis invariance, called Maximal Axis Projection (MAP), that adopts the permutation-invariant axis projection functions to determine the canonical directions.
+Unfortunately, when the eigenvalues separation is small due to symmetries in the geometries, these methods becomes unstable and sensitive to noise.
+To overcome this issue, in this work we propose the Coupled Laplacian operator to produce a-priori aligned eigenmaps for several registered shapes, and perform locally-aware shape matching.
+Deep-learned point-cloud descriptors [24, 72, 6, 77] arose as an alternative to generate local features for 3D sufraces.
+We can distinguish three main categories depending on the backbone architecture employed.
 Convolutional Neural Networks (CNNs) are often used on point clouds projected into 2D depth images [16, 29] or directly on 3D voxels [24, 77, 29].
-For example, 3DMatch [77] aims to learn a local volumetric patch descriptor for establishing correspondences between partial 3D data.
 Secondly, to work directly on raw point cloud data, PointNet and PointNet++ have been proposed aiming to learn rotation and permutation invariant features [49, 50].
-Based on PointNet, several learned descriptors have been proposed [15, 14, 75, 76].
+Based on PointNet, several learned descriptors have been introduced [14, 13, 75, 76].
 Among them, PPFNet [14] and PPF-FoldNet [13] try to improve the feature representation of PointNet by incorporating global context and point-pair features.
@@
 Furthermore, in supervised algorithms [77, 24, 14, 13], paired 3D patches, such as in the 3DMatch dataset [77], are needed to train the models.
-For this reason, the 3D rigid point cloud matching method we propose here does not need training and utilizes only the similarity among properly aligned spectral embeddings to relate an unseen target shape to the given references.
+For this reason, the 3D rigid point cloud matching method we propose does not need training and utilizes only the similarity among properly aligned spectral embeddings to relate an unseen target shape to the given references.
 3 Method
@@
 where $\boldsymbol{B}\in\mathbb{R}^{n\times n}$ is generally set as $\boldsymbol{D}$, and $\lambda_{i}\leq\lambda_{i+1}$.
-Finally, eigenmaps are simply eigenvectors sub-spaces, generated by leaving out $\boldsymbol{\phi}_{0}$, corresponding to $\lambda_{0}=0$ and using the next $m$ eigenvectors for embedding graph nodes in an $m$-dimensional space, $\boldsymbol{x}_{i}\rightarrow[\boldsymbol{\phi}_{1}(i),\cdots,\boldsymbol{\phi}_{m}(i)]$
+Finally, eigenmaps are simply eigenvectors sub-spaces, generated by leaving out $\boldsymbol{\phi}_{0}$, corresponding to $\lambda_{0}=0$, and using the next $m$ eigenvectors for embedding graph nodes in an $m$-dimensional space, $\boldsymbol{x}_{i}\rightarrow[\boldsymbol{\phi}_{1}(i),\cdots,\boldsymbol{\phi}_{m}(i)]$
 Coupled Laplacian.
@@
 2.
-Let $\mathcal{G}^{\mathcal{T}}$ be the graph corresponding to a target 3D point cloud and $\mathcal{G}^{S_{1}},\cdots,\mathcal{G}^{S_{N}}$ the ones obtained from $N$ different sources.
+Let $\mathcal{G}^{\mathcal{T}}$ be the graph of a target 3D point cloud and $\mathcal{G}^{S_{1}},\cdots,\mathcal{G}^{S_{N}}$ the ones of $N$ different sources.
 We construct a global graph, or coupled graph, $\mathcal{G}^{C}$, by adding cross-connections separately between the vertices of the source, $\mathcal{V}^{\mathcal{T}}=\{\boldsymbol{x}^{\mathcal{T}}_{i}\}_{i=1}^{n^{\mathcal{T}}}$, and each of the reference shapes, $\mathcal{V}^{S_{k}}=\{\boldsymbol{x}^{S_{k}}_{i}\}_{i=1}^{n^{S_{k}}}$ with $k=1,\cdots,N$.
-To include meaningful cross-connections, we first perform a rigid [20, 21, 45] or affine [45] registration of the source geometries to better match the target.
-The latters are preferred to a non-rigid registration because, when dealing with the identification of fine local variations, they do not change the relative position of the points inside a point-set and, therefore, the deformations are kept almost unchanged.
-After that, a sub-set of vertices, $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ , is stochastically extracted from the target geometry and its nearest correspondence is searched in each of the aligned reference point clouds, such as
-$F^{S_{k}}=\{\boldsymbol{f}^{S_{k}}_{j}:\min_{\boldsymbol{x}^{S_{k}}_{i}\in\mathcal{V}^{S_{k}}}\text{d}^{2}(\boldsymbol{x}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{j})|\boldsymbol{f}^{\mathcal{T}}_{j}\in F^{\mathcal{T}}\}$ (3)
+To include meaningful cross-connections, we first perform a rigid [19, 20, 45] or affine [45] registration of the source geometries to align to the target.
+These methods are preferred to a non-rigid registration because, when dealing with the identification of fine local variations, they do not change the relative position of the points inside a point-set and, therefore, the deformations are kept unchanged.
+After that, a sub-set of vertices, $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ , is stochastically extracted from the target geometry and their nearest correspondences are searched in each of the aligned reference point clouds as
+$F^{S_{k}}=\{\boldsymbol{f}^{S_{k}}_{j}:\operatorname*{arg\,min}_{\boldsymbol{x}^{S_{k}}_{i}\in\mathcal{V}^{S_{k}}}\text{d}^{2}(\boldsymbol{x}^{S_{k}}_{i},\boldsymbol{f}^{\mathcal{T}}_{j})|\boldsymbol{f}^{\mathcal{T}}_{j}\in F^{\mathcal{T}}\}$ (3)
 with $k=1,\cdots,N$.
-In this way, there is no constraint on the original number points comprising each shape nor on their initial coordinates system, e.g.
-camera coordinates.
+In this way, there is no constraint on the original number points comprising each shape nor on their initial coordinates systems.
 Furthermore, the cardinality of the target sub-set, is chosen as a fraction of its total points, $|F^{\mathcal{T}}|=l\cdot n^{\mathcal{T}}$, where $0<l\leq 1$.
@@
 Thanks to the vertex ordering of the coupled graph, the solution of the generalized eigenproblem of Eq.
-2 applied to the Coupled Laplacian with $\boldsymbol{B}^{C}=\text{diag}([\boldsymbol{D}^{\mathcal{T}},\boldsymbol{D}^{S_{1}},\cdots,\boldsymbol{D}^{S_{N}}])$, which is the stack of uncoupled diagonal degree matrices, gives coupled eigenvectors $\{\boldsymbol{\phi}^{C}_{i}\}_{i=0}^{n}$.
+2 applied to the Coupled Laplacian, with $\boldsymbol{B}^{C}=\text{diag}([\boldsymbol{D}^{\mathcal{T}},\boldsymbol{D}^{S_{1}},\cdots,\boldsymbol{D}^{S_{N}}])$, yields to coupled eigenvectors $\{\boldsymbol{\phi}^{C}_{i}\}_{i=0}^{n}$.
 These eigenvectors can then be split into each single component of the coupled graph as
@@
 The first one can be obtained by measuring the similarity between the aligned eigenspaces through the Grassmann distance, $d_{G}(\cdot,\cdot)$ [74].
-Nevertheless, it cannot be computed directly, since we consider point-sets of arbitrary size and we do not have point-to-point correspondences.
+Nevertheless, it cannot be computed directly, since we consider point clouds of arbitrary size and we do not have point-to-point correspondences.
 Hence, we restrict the distance computation only to the set of cross-connected vertices, $F^{\mathcal{T}}$ for the target and $F^{S_{k}}$ for the $k$-th reference.
 In order to get the reduced basis restricted only to those coupled points, we perform a QR decomposition of the restriction of the eigenmodes to the cross-points.
-The best matching source is therefore given by
+The best matching source is then given by
 $\operatorname*{arg\,min}_{k}d_{G}(\boldsymbol{Q}^{\mathcal{T}},\boldsymbol{Q}^{S_{k}}).$ (5)
@@
 This first approach is used in the following to solve the BSE task by using two source shapes, i.e.
-the reference bone of known side and its mirrored version.
+the reference bone of known side and its contralateral mirrored version.
 On the other hand, to obtain point-wise similarity scores we propose to compare the $m$-dimensional embeddings of the cross-connection vertices by using the cosine distance function.
@@
 Dataset Femur Hip Tibia Fibula
-L
-R
-L
-R
-L
-R
-L
-R
-Human
-Fisher et al. [19]
-18
-19
-20
-20
-SSM-Tibia [32]
-30
-30
-ICL [47]
-35
-35
-35
-35
-35
-35
+L R L R L R L R
+Human Fisher et al. [18] 18 19 20 20 - - - -
+SSM-Tibia [31] - - - - - 30 - 30
+ICL [47] 35 35 - - 35 35 35 35
 S
-Internal [40]
-18
-18
-18
-18
-Total
-71
-72
-20
-20
-53
-83
-35
-65
+Internal [40] 18 18 - - 18 18 - -
+Total 71 72 20 20 53 83 35 65
 Table 2: Anomaly localization results.
@@
 GAN, AE and VM results are provided by Bergmann et al.
-[7], and, for a fair comparison, we include the results obtained by restricted GPS [55], eigenfunctions (Hist) [42, 60] and Euclidean matching.
-The subsctipt on a method indicates the number of used eigenmaps.
+[7].
+Moreover, we include the results obtained by restricted GPS [55], eigenfunctions (Hist) [42, 60] and Euclidean matching and MAP [39].
+All the matching methods are applied after affine CPD registration, when not specified, or CMM + CPD Non-Rigid (NR) registration.
+The subsctipt on a method indicates the number of eigenmaps used.
 The overall best performing methods are highlighted in boldface, while the bests for each category are underlined.
-Method
-Bagel
-Cable Gland
-Carrot
-Cookie
-Dowel
-Foam
-Peach
-Potato
-Rope
-Tire
-Mean
-3D RGB
-VM
-0.388
-0.321
-0.194
-0.570
-0.408
-0.282
-0.244
-0.349
-0.268
-0.331
-0.335
-GAN
-0.421
-0.422
-0.778
-0.696
-0.494
-0.252
-0.285 0.362
-0.402
-0.631
-0.474
-AE 0.432
-0.158
-0.808
-0.491
-0.841 0.406
-0.262
-0.216
-0.716
-0.478
-0.481
-3D Only
-GAN
-0.111
-0.072
-0.212
-0.174
-0.160
-0.128
-0.003
-0.042
-0.446
-0.075
-0.143
-AE
-0.147
-0.069
-0.293
-0.217
-0.207
-0.181
-0.164
-0.066
-0.545
-0.142
-0.203
-VM
-0.280
-0.374
-0.243
-0.526
-0.485
-0.314
-0.199
-0.388
-0.543
-0.385
-0.374
-[…]
-0.452
-0.616
-0.695
-0.364
-0.738
-0.471
-0.659
-0.844
-0.647
-0.651
-0.613
-[…]
-0.465
-0.621
-0.690
-0.363
-0.739
-0.480
-0.672
-0.833
-0.653
-0.670
-0.619
-[…]
-0.476
-0.629
-0.703
-0.365
-0.744
-0.473
-0.661
-0.840
-0.647
-0.693
-0.623
-[…]
-0.491
-0.629
-0.698
-0.351
-0.746
-0.501
-0.663
-0.841
-0.652
-0.695
-0.627
-Euclidean
-0.655
-0.631
-0.743
-0.615
-0.803
-0.528
-0.726
-0.875
-0.762
-0.695
-0.703
-$\textbf{Ours}_{100}$
-0.669
-0.642 0.808
-0.714
-0.812
-0.582
-0.748
-0.897
-0.750
-0.733
-0.736
-$\textbf{Ours}_{200}$ 0.702
-0.630
-0.728
-0.735 0.812 0.701 0.780 0.914 0.767
-0.713
-0.748
-Table 3: Accuracy of human BSE for each bone structures. All the matching methods are applied after RANSAC registration. The overall best performing methods are highlighted in boldface.
-Method
-Femur
-Hip
-Fibula
-Tibia
-Mean
-Hausdorff
-73.52
-97.95
-59.30
-72.52
-75.82
-Chamfer
-71.64
-98.65
-64.62
-74.47
-77.35
-FPFH [57]
-68.88
-96.67
-66.69
-78.35
-77.65
-$\textbf{Ours}_{20}$
-78.79
-98.78
-71.28
-78.46
-81.83
-$\textbf{Ours}_{10}$ 79.68
-97.76
-73.47 78.66 82.39
+Method Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Mean $\uparrow$
+3D RGB VM 0.388 0.321 0.194 0.570 0.408 0.282 0.244 0.349 0.268 0.331 0.335
+GAN 0.421 0.422 0.778 0.696 0.494 0.252 0.285 0.362 0.402 0.631 0.474
+AE 0.432 0.158 0.808 0.491 0.841 0.406 0.262 0.216 0.716 0.478 0.481
+3D Only GAN 0.111 0.072 0.212 0.174 0.160 0.128 0.003 0.042 0.446 0.075 0.143
+AE 0.147 0.069 0.293 0.217 0.207 0.181 0.164 0.066 0.545 0.142 0.203
+VM 0.280 0.374 0.243 0.526 0.485 0.314 0.199 0.388 0.543 0.385 0.374
+Euclidean (NR) 0.404 0.623 0.731 0.366 0.771 0.303 0.590 0.772 0.697 0.583 0.584
+GPS100[55] 0.452 0.616 0.695 0.364 0.738 0.471 0.659 0.844 0.647 0.651 0.613
+GPS200[55] 0.465 0.621 0.690 0.363 0.739 0.480 0.672 0.833 0.653 0.670 0.619
+Hist100 [60] 0.476 0.629 0.703 0.365 0.744 0.473 0.661 0.840 0.647 0.693 0.623
+MAP200 [39] 0.481 0.630 0.694 0.399 0.742 0.497 0.653 0.832 0.649 0.675 0.625
+Hist200 [60] 0.491 0.629 0.698 0.351 0.746 0.501 0.663 0.841 0.652 0.695 0.627
+Euclidean 0.655 0.631 0.743 0.615 0.803 0.528 0.726 0.875 0.762 0.695 0.703
+$\textbf{Ours}_{100}$ 0.669 0.642 0.808 0.714 0.812 0.582 0.748 0.897 0.750 0.733 0.736
+$\textbf{Ours}_{200}$ 0.702 0.630 0.728 0.735 0.812 0.701 0.780 0.914 0.767 0.713 0.748
+Table 3: Accuracy [%] of human BSE.
+All the matching methods are applied after RANSAC registration with spectral scaling, when not specified, or CMM + CPD Non-Rigid (NR) registration.
+The overall best performing methods are highlighted in boldface.
+Method Femur Hip Fibula Tibia Mean $\uparrow$
+Chamfer (NR) 56.58 86.58 57.89 74.74 68.95
+Hausdorff (NR) 57.37 87.63 59.47 73.68 69.54
+Hausdorff 73.52 97.95 59.30 72.52 75.82
+Chamfer 71.64 98.65 64.62 74.47 77.35
+FPFH [57] 68.88 96.67 66.69 78.35 77.65
+MAP20 [39] 76.05 98.54 71.32 74.21 80.13
+Hist20 [60] 77.63 98.42 69.47 77.37 80.72
+$\textbf{Ours}_{20}$ 78.79 98.78 71.28 78.46 81.83
+$\textbf{Ours}_{10}$ 79.68 97.76 73.47 78.66 82.39
 Bone Side Estimation.
-The proposed global shape matching score is used to perform automatic BSE, that is, given a reference bone of known side we want to assess the side of another bone, referred as target.
-Assuming general unknown initial frames, we used the Principal Component Analysis (PCA) to generate a mirrored version of the reference surface.
+The proposed global shape matching score is used to perform automatic BSE.
+Assuming general unknown initial frames, we used the Principal Component Analysis (PCA) to generate a mirrored version of the reference surface, i.e.
+a synthetic contralateral bone.
 Namely, we identified that the bilateral symmetry of human and animal bodies is equivalent to a mirroring around the second principal component on the vast majority of bones, such as the ones considered in this study.
@@
 2 can be applied.
-More specifically, we selected the Random Sample Consensus (RANSAC) [20], preceded by a spectral scaling, as rigid registration method.
+More specifically, we selected the Random Sample Consensus (RANSAC) [19], preceded by a spectral scaling, as rigid registration method to handle varying bone lengths without altering local structures.
 The scaling is performed using only the information carried by the Fiedeler vector, i.e.
@@
 Note that, the proposed BSE method is fully independent of the target and reference frames and therefore can be applied directly on the segmentation obtained from a medical image, without any previous knowledge, in a fully automated pipeline.
-Additionally, in alternative, one can using two different bone references, one left and one right, avoiding the mirroring step, or apply the mirroring on the target shape while letting the source unchanged.
+In alternative, one can using two different bone references, one left and one right, avoiding the mirroring step, or apply the mirroring on the target shape while letting the source unchanged.
 Full details of the described algorithm are provided in the Supplementary Material.
@@
 [18] we extracted Femur and Hip structures, from SSM-Tibia [31] only right Tibia and Fibula, and from the Imperial College London (ICL) [47] we collected Femur, Tibia and Fibula point clouds.
-The composition of the benchmark is shown in Tab.
-1 and the detailed instructions to collect and prepare it for testing are provided in the Supplementary Code (or GitHub).
+The detailed composition of the benchmark is reported in Tab.
+1.
 Such a variety of shapes, which are acquired using different methods, makes the task more challenging, including some intra-clinic variability.
-Some additional internal data [40] with knee structures of sheep are used in the Supplementary Material to test the human to animal generalization, and vice-versa, of the proposed BSE technique.
-In our experiments, for a given bone class, we selected each shape once as source and the others as targets on which the side has to be inferred.
-Therefore, the average accuracy of all the experiments takes into account the variability of the source shape, and the method robustness can be analysed.
+In our experiments, for a given bone class, we performed cross-testing by selecting each shape once as source and the others as targets on which the side has to be inferred.
+Therefore, the average accuracy of all the experiments, achieved when predicting the correct body side, i.e.
+binary classification, takes into account the variability of the source shape.
+The robustness of the method with respect to the reference is discussed in the Supplementary Material.
 Figure 3: Average BSE accuracy with respect to the fraction of cross-edges used to build the Coupled Laplacian and the number of eigenmaps used for matching.
@@
 In our experiments, we chose one anomaly-free training sample for each class as source and we used the point-wise distances of the aligned eigenmaps between this shape and each target in the test set as anomaly scores.
-In this case, we performed the registration using affine Coherent Point Drift (CPD) [45], which is more suited for objects like the ones included in the MVTec 3D-AD, e.g.
-the length and thickness variation among potatoes and carrots samples is better captured by an affine registration rather than rigid.
-In order to speed-up the computation, we pre-processed the point clouds by removing their flat background with a threshold on the $z$-axis after a 3-dimensional PCA and, without losing generality, we sub-sampled the number of foreground points to a maximum of $13000$.
-Finally, to compare the obtained anomaly scores with the Ground Truth (GT) image, we projected back the points to the original 2D plane and, if any sub-sampling was performed in pre-processing, we applied a dilation with a structuring element of size the inverse of the sampling factor.
+We performed the pre-registration using affine Coherent Point Drift (CPD) [45], which is more suited for objects like the ones included in the MVTec 3D-AD, e.g.
+length and thickness variations are better captured by an affine registration rather than rigid.
+In order to speed-up the computation, we pre-processed the point clouds by removing their flat backgrounds with a threshold on the $z$-axis after a 3-dimensional PCA and, without losing generality, we sub-sampled the number of foreground points to a maximum of $13000$.
+Finally, to compare the obtained anomaly scores with the Ground Truth (GT) image, we projected back the points to the original 2D plane and, if any sub-sampling was performed, we applied a dilation with a structuring element of size the inverse of the sampling factor.
 4.2 Experiment results
-We show here comparative quantitative and qualitative results of the proposed method, while a study of the robustness with respect to the source shape and more graphical comparisons are provided in the Supplementary Material.
 Bone Side Estimation.
 In Fig.
-3, the average accuracy of BSE is depicted based on the fraction of artificial cross-edges added to compute the Coupled Laplacian and the dimension of the spectral embeddings, $l$ and $m$, respectively.
+3, the average accuracy of BSE is depicted based on the fraction of artificial cross-edges added to create the coupled graph and the dimension of the spectral embeddings, $l$ and $m$, respectively.
 We can observe that the more coupled eigenmaps we aim to utilize for computing a global similarity score, the greater the need for additional cross-connections to ensure the reliability of the results.
 This occurs because, when graphs are weakly coupled with only a few cross-connections, the Coupled Laplacian captures limited intra-shape characteristics and predominantly emphasizes individual local geometries.
-In fact, when $l=0$ and therefore the coupled graph is not fully connected, the eigenmaps obtained using the Coupled Laplacian are equivalent to the ones computed independently on each single graph, if the eigenvalues are the same (see Supplmentary Material).
+In fact, when $l=0$ and therefore the coupled graph is not fully connected, the eigenmaps obtained using the Coupled Laplacian are equivalent to the ones computed independently on each single graph, if the eigenvalues are the same (proof in the Supplmentary Material).
 On the other hand, not too many cross-edges, nor eigenmaps, are needed to obtain meaningful global matching scores and achieve good performances in the BSE task.
 In Tab.
-3 a quantitative comparison of human BSE with respect to other distance measures is reported.
-For a fair compariosn, all the other metrics are computed after RANSAC registration of source and mirrored source.
-Hausdorff and Chamfer discrepancies are calculated on the Euclidean coordinates of the points, while FPFH [57] are 33-dimensional vectors describing each point based on histograms of the surface normals.
-Our method, using $10$ and $20$ eigenmodes with $l=0.5$ fraction of cross-connections, outperforms the other techniques achieving better accuracy both on the single bones and in average.
-This is because the description provided by the aligned eigenmaps, obtained through Coupled Laplacian, is more aware of local details than the other distance metrics.
+3 a quantitative comparison with other methods of human BSE accuracy is reported.
+For a fair comparison, all methods are performed after RANSAC registration of the sources.
+Hausdorff and Chamfer discrepancies are calculated on the Euclidean coordinates of the points, for completeness, also in case of 2-Step non-rigid GMM + CPD registration ($2000$x slower than RANSAC).
+Our method, using $10$ and $20$ eigenmodes with $l=0.5$ fraction of cross-connections, outperforms the other techniques achieving higher accuracy both on the single bones and in average.
+Hence, the description provided by the aligned eigenmaps, obtained with Coupled Laplacian, is more aware of local details than the other methods.
 Anomaly Localization.
 Tab.
-2 lists quantitative results for each evaluated method for the localization of anomalies.
-For each category, the normalized area under the Per-Region Overlap (PRO) curve [7] with an upper integration limit of $0.3$, as well as, the mean performance, are reported.
+2 lists quantitative results of each evaluated method for the localization of anomalies.
+For each category, the normalized area under the Per-Region Overlap (PRO) curve with an upper integration limit of $0.3$ [7], as well as, the mean performance, are reported.
 Performance of Generative Adversarial Network (GAN), Autoencoder (AE) and Variation Model (VM) on the same test set are provided by the dataset authors [7].
-Furthermore, we tested restricted GPS [55], eigenmaps histogram matching [42, 60] and our method, all with the same pre-processing and source shapes, using $100$ and $200$ eigenmaps.
-In order to obtain denser and more precise anomaly localization, the Coupled Laplacian is built using the whole set of points, after pre-processing, as cross-connections, i.e.
+Furthermore, we tested restricted GPS [55], eigenmaps histogram matching [42, 60], MAP [39] and our method, all with the same pre-processing and source shapes, using $100$ and $200$ eigenmaps.
+In order to obtain denser and more precise anomaly localization, the coupled graph is built using the whole set of points as cross-connections, i.e.
 $l=1$.
-Furthermore, we include the result obtained using as anomaly score for each target point the normalized Euclidean distance of the nearest point in the source geometry after registration.
-Our method, using only 3D information, outperforms all the other techniques in average and on most of the individual objects.
+Furthermore, we include the result obtained using as anomaly score for each target point the normalized Euclidean distance of the nearest point in the source geometry, both using rigid and non-rigid registration.
+Our method, using only 3D information, outperforms all the other techniques.
 Moreover, we obtained better results than deep-learning methods having RGB and Depth in combination as input.
@@
 This concept is linked to the modal length, which we define in the Supplementary Material.
-The trade-off between number of maps used and extent of surface differences detected is interesting to tune the proposed method to other tasks requiring specific attention to identify surface dissimilarities.
-Figure 5: Comparison of the distance between source and target, in the 3D anomaly detection task, using the Coupled Laplacian technique and Euclidean distance after RANSAC rigid alignment.
+The trade-off between number of maps and extent of differences detected is interesting to tune the method to other tasks requiring specific attention to identify surface dissimilarities.
+Figure 5: Comparison of the anomaly score obtained using the Coupled Laplacian technique, with $m=200$ and $l=1$, and Euclidean distance, both after affine CPD alignment.
 Ablation Study.
-To further motivate the use of the Laplacian spectral space over the Euclidean space, in Fig.
-5 we show some comparison of the distance between source and target, after RANSAC rigid alignment, in the two spaces, the latter obtained using the proposed Coupled Laplacian operator with $m=200$.
+To further motivate the preference for Coupled Laplacian spectral space over the Euclidean one, in Fig.
+5, we display comparisons of point-wise distances (anomaly scores) between source and target in both spaces, after the same registration.
 In the first and second row the better resistance to noise and outliers in the spectral space is highlighted.
 The bending of the cable gland, i.e.
-the anomaly, is not detected in the Euclidean space because of an additional segment of points on the back of the target having the largest distance to the source.
-On the other hand, with the aligned eigenmaps comparison, the anomaly, as well as the noise, is correctly localized.
-This is because both are local differences, with respect to reference, and, therefore, they can be easily captured in the Laplacian space.
+the anomaly, is barely detected in the Euclidean space because of noisy points on the back.
+Instead, with the aligned eigenmaps comparison, the anomaly, as well as the noise, is correctly localized.
+This is because both are local differences, with respect to the reference, and, therefore, they can be easily captured with coupled eigenmaps.
 In addition, in the third row, we show that our technique provides a sharper anomaly split of the cookie point cloud, which simplifies the anomaly detection task.
@@
 Moreover, a wrong registration of the shapes, due to randomness and errors in the selected registration algorithm, may affect the alignment of the $m$-dimensional embeddings derived from the Coupled Laplacian.
-Even if the probability of such event for similar geometries is relatively low and the adjacency matrix is already weighted with RBF, a double-check can be added in the pipeline to verify the correct registration [13].
+We report here some statistics on registration failures in BSE (failure rate [%] / BSE accuracy on failures [%]), using different references, for femur 6.84 $\pm$ 7.82 / 96.67 $\pm$ 11.30 (high symmetry) and hip 0.0 $\pm$ 0.0 (low symmetry).
+Moreover, to test robustness, we added white Gaussian rotation and translation noise, i.i.d.
+in the 3 axis post-registration, obtaining (Rot.
+SD [] / Trans.
+SD [mm] / Acc.
+Drop [%]): 2.5 / 5 / 8.16 - 5 / 5 / 15.13 - 7.5 / 10 / 18.16.
+Given the low registration failure rates and high robustness to noise, we can rely on the proposed coupling algorithm, making the method independent of the acquisition frames.
+Moreover, the method can well generalize to any chiral shape matching problem and to tasks like identification of topological noise in graphs and partial matching (see example in Supplementary Material).
 5 Conclusions
@@
 Leveraging global similarities of eigenmaps derived from the Coupled Laplacian, we define a full pipeline to perform BSE on arbitrary bone surfaces segmented from a generic medical image.
-Moreover, we tested the ability of the proposed method in capturing surface differences by performing local shape matching for the 3D anomaly detection task on the MVTec 3D-AD dataset.
+Moreover, we tested the ability of the proposed method in capturing local surface differences by performing 3D anomaly detection on the MVTec 3D-AD dataset.
 The proposed technique outperforms other methods on the two tasks, therefore capturing better both global and local similarities, thanks to the matching performed on correctly aligned spectral spaces.
@@
 $\alpha\leftarrow L^{S}/L^{\mathcal{T}}$ ;
-  normal-$\triangleright$ Scale factor
+  $\triangleright$ Scale factor
 $\mathcal{V}^{\mathcal{T}}\leftarrow\alpha\times\mathcal{V}^{\mathcal{T}}$;
@@
 $\mathcal{V}^{S},\mathcal{V}^{S}_{M}\leftarrow$ RANSAC registrations to $\mathcal{V}^{\mathcal{T}}$;
-Select target cross-connections set $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ Find sets $F^{\mathcal{S}}\subset\mathcal{V}^{S}$ and $F^{\mathcal{S}}_{M}\subset\mathcal{V}^{S}_{M}$ as in Eq.
-3 $L^{C}\leftarrow\text{Coupled Laplacin}(W^{\mathcal{T}},W^{S},F^{\mathcal{T}},F^{\mathcal{S}},F^{\mathcal{S}}_{M})$;
+Select target cross-connections set $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$
+Find sets $F^{\mathcal{S}}\subset\mathcal{V}^{S}$ and $F^{\mathcal{S}}_{M}\subset\mathcal{V}^{S}_{M}$ as in Eq. 3
+$L^{C}\leftarrow\text{Coupled Laplacin}(W^{\mathcal{T}},W^{S},F^{\mathcal{T}},F^{\mathcal{S}},F^{\mathcal{S}}_{M})$;
 $\{\lambda_{i}^{C}\}_{i=0}^{m}$, $\{\boldsymbol{\phi}_{i}^{C}\}_{i=0}^{m}\leftarrow$ Eig$(L^{C},m)$;
@@
       $\mathrm{S}^{\mathcal{T}}$ = $\mathrm{S}^{S}$ ;
-        normal-$\triangleright$ Same side
+        $\triangleright$ Same side
 else
       $\mathrm{S}^{\mathcal{T}}$ = $\neg\mathrm{S}^{S}$ ;
-        normal-$\triangleright$ Opposite side
+        $\triangleright$ Opposite side
 end if
@@
 The ratio between source and target lengths gives a scaling factor, $\alpha$, that we use to scale the target surface in order to match the length of the reference.
-The mirrored version of the source bone, $\mathcal{V}^{S}_{M}$, is then generated by flipping it around the point cloud second principal component and both versions of the reference are registered to the target with RANSAC algorithm [20].
+The mirrored version of the source bone, $\mathcal{V}^{S}_{M}$, representing its contralateral, is then generated by flipping it around the point cloud second principal component and both versions of the reference are registered to the target with RANSAC algorithm [19].
 Interestingly, the adjacency matrix of the two references, before and after registration, is the same as the original one and, therefore, we can avoid its computation multiple times.
@@
 3.
+Alternatively, assuming not flat surfaces, other methods for cross-connections can be used instead of NN, e.g.
+spectral clustering medoids.
 Hence, the three graphs can now be connected and the Coupled Laplacian computed from the whole weighted adjacency matrix.
@@
 The overall best performing methods are highlighted in boldface.
-Method
-Femur
-Hip
-Fibula
-Tibia
-Best Hausdorff
-89.62
-100.0
-75.76
-86.87
-Chamfer
-85.85
-100.0
-86.87
-90.91
-FPFH [57]
-83.96
-100.0
-83.84
-93.94
-$\textbf{Ours}_{20}$ 93.40 100.0
-89.90
-94.95
-$\textbf{Ours}_{10}$ 93.40 100.0 91.92
-93.94
-Worst Hausdorff
-51.89
-89.74
-39.39
-51.51
-Chamfer
-54.72
-92.31
-35.35
-38.38
-FPFH [57]
-34.34
-87.18
-34.34
-41.41
-$\textbf{Ours}_{20}$
-49.06
-94.87
-38.38
-39.39
-$\textbf{Ours}_{10}$ 57.55
-84.62
-43.43
-45.45
+Method Femur Hip Fibula Tibia
+Best Hausdorff 89.62 100.0 75.76 86.87
+Chamfer 85.85 100.0 86.87 90.91
+FPFH [57] 83.96 100.0 83.84 93.94
+$\textbf{Ours}_{20}$ 93.40 100.0 89.90 94.95
+$\textbf{Ours}_{10}$ 93.40 100.0 91.92 93.94
+Worst Hausdorff 51.89 89.74 39.39 51.51
+Chamfer 54.72 92.31 35.35 38.38
+FPFH [57] 34.34 87.18 34.34 41.41
+$\textbf{Ours}_{20}$ 49.06 94.87 38.38 39.39
+$\textbf{Ours}_{10}$ 57.55 84.62 43.43 45.45
 Bone Side Estimation.
@@
 S and H stands for Sheep and Human, respectively, and the overall best performing methods are highlighted in boldface.
-Method
-Femur
-Tibia
-Mean
-S normal-$\rightarrow$ S Hausdorff
-94.37
-97.14
-90.48
-Chamfer
-94.29
-88.57
-94.88
-FPFH [57]
-97.14
-97.14
-97.14
+Method Femur Tibia Mean
+S $\rightarrow$ S Hausdorff 94.37 97.14 90.48
+Chamfer 94.29 88.57 94.88
+FPFH [57] 97.14 97.14 97.14
 $\textbf{Ours}_{20}$ 97.30 97.30 97.30
 $\textbf{Ours}_{10}$ 97.30 97.30 97.30
-H normal-$\rightarrow$ S Hausdorff
-72.22
-69.44
-70.83
-Chamfer
-69.44
-75.00
-72.22
-FPFH [57]
-77.78
-75.00
-76.39
-$\textbf{Ours}_{20}$
-69.44
-72.22
-70.83
+H $\rightarrow$ S Hausdorff 72.22 69.44 70.83
+Chamfer 69.44 75.00 72.22
+FPFH [57] 77.78 75.00 76.39
+$\textbf{Ours}_{20}$ 69.44 72.22 70.83
 $\textbf{Ours}_{10}$ 83.33 77.78 80.56
-S normal-$\rightarrow$ H Hausdorff
-57.94
-63.00
-60.47
-Chamfer
-62.62
-66.00
-64.31
-FPFH [57] 63.89
-63.00
-63.45
-$\textbf{Ours}_{20}$
-63.55
-60.00
-61.78
-$\textbf{Ours}_{10}$
-63.55
-66.00 64.78
+S $\rightarrow$ H Hausdorff 57.94 63.00 60.47
+Chamfer 62.62 66.00 64.31
+FPFH [57] 63.89 63.00 63.45
+$\textbf{Ours}_{20}$ 63.55 60.00 61.78
+$\textbf{Ours}_{10}$ 63.55 66.00 64.78
 Figure 14: Comparison of human and sheep femur and tibia bones. For both of them, the top and lower parts are shown separately to highlight their differences.
@@
 Addressing this challenge will contribute to advancing the field of BSE and refining the understanding of global similarities and localized distinctions in diverse bone structures.
+12 Scalability
+Figure 15: Average computational time of the eigendecomposition using different combinations of number of shapes ($1$, $3$ or $5$) and modes ($10$ or $20$).
+Figure 16: Coupled modes correctly aligned in a partial matching scenario.
+The efficiency trade-off of decomposing larger Laplacians is illustrated in Fig.
+15.
+The average computational time is reported for the eigendecomposition, using $10$ and $20$ maps, coupling the same shape with itself $3$ or $5$ times.
+As expected, larger, i.e.
+number of shapes, and denser, i.e.
+number of cross-connections, is the Laplacian, higher is the time needed for the computation.
+Scalability might be a concern in tasks like shape retrieval with numerous classes.
+However, simple measures, e.g.
+downsampling shapes and interpolating eigenmaps, can reduce the complexity.
+Given the intrinsic coupling of modes, faster, but randomized, solvers, e.g.
+AMG, are still precise and have been used, cutting the need of resources.
+13 Partial Matching
+In Fig.
+16 we show a graphical example of coupled eigenmaps derived from partially matching shapes.
+More in detail, an entire femur bone is matched with the distal part, i.e.
+lower part, of another femur.
+In this case, artificial cross-connections are added only on the partial match of the source shapes, leading to correctly aligned eigenmaps.
+Namely, the embeddings are successfully coupled in the distal area of the femur while they are independent on the remaining surface of the source.
+This example shows the potential of the Coupled Laplacian operator in different, and more complex, scenarios.